### v2

# Mixed Finite element method for stress gradient elasticity

Abstract.
This paper develops stable finite element pairs for the linear stress gradient elasticity model, overcoming classical elasticity’s limitations in capturing size effects. We analyze mesh conditions to establish parameter-robust error estimates for the proposed pairs, achieving unconditional stability for finite elements with higher vertex continuity and conditional stability for CG-DG pairs when no interior vertex has edges lying on three or fewer lines. Numerical experiments validate the theoretical results, demonstrating optimal convergence rates.
The work of TL was supported by the NSFC Project 123B2014. The work of ST was supported by the NSFC 12401483.
1. Introduction
Classical linear elasticity models have been widely applied to fundamental problems and engineering applications. As a consequence, numerous structure-preserving and efficient finite element methods for these models are proposed and well-developed in literature, such as Arnold2002 ; Arnoldweaksym ; Arnold3D ; HuZhang14 ; HZ3D ; HZfamily ; HZlowest ; GuzmanAfiled ; PaulyZulehner23 . Nevertheless, for materials with microstructure, the classical elasticity model lacks an internal length scale in its constitutive structure, making it unable to capture phenomena such as size effects Aifantis2009 ; Aifantis2011overview . Consequently, classical continuum theories fall short in providing an accurate and detailed description of deformation phenomena in such materials observed in experiments Aifantis2011overview . This motivates further exploration of the generalization either by introducing modifying terms or by allowing certain laws to be violated.
Gradient elasticity theories MINDLIN1964 ; Aifantis1986 extend classical elasticity by incorporating higher-order spatial derivatives of strain, stress, etc. The introduction of higher-order terms of strain or stress can handle the singularity caused by the crack tips and dislocation cores Aifantis2011overview . In this article, we focus on the linear stress gradient elasticity model, first introduced by Eringen Eringen1983 , as a model problem of the gradient elasticity theory. In the setting of stress gradient theory, the constitution law is replaced by $\mathcal{E}=\mathcal{A}(\sigma-\iota^{2}\Delta\sigma)$, where the size parameter $\iota$ is a positive constant and we assue that $\iota\leq 1$ in this paper. This problem has been discussed in engineering Aifantis2011overview ; VariationLSG , but there are few relevant theoretical results in numerical analysis so far.
Given a two-dimensional connected domain $\Omega$, the linear stress gradient elasticity equation is
(1.1) $\begin{cases}\mathcal{E}=\mathcal{A}(\sigma-\iota^{2}\triangle\sigma)&\text{in }\Omega,\\ \mathrm{div}\sigma=f&\text{in }\Omega,\\ 2\mathcal{E}=\mathrm{symgrad}~u&\text{in }\Omega,\\ \sigma n=g_{f},~&\text{on }\Gamma_{f},\\ u=g_{c},~&\text{on }\Gamma_{c},\\ \partial_{n}\sigma=0,&\text{on }\partial\Omega.\end{cases}$
where $\Gamma_{c}$ and $\Gamma_{f}$ is a partition of $\partial\Omega$. Here $\sigma$, $\mathcal{E}$, and $u$ represent the stress (symmetric matrix-valued), strain (symmetric matrix-valued), and displacement (vector-valued), respectively. In this paper, we assume $\mathcal{A}$ is an isotropic fourth-order tensor, defined as
$\mathcal{A}\sigma=\frac{1}{2\mu}\sigma-\frac{\lambda}{2\mu(2\mu+2\lambda)}\mathrm{tr}(\sigma)I.$
Clearly, $\mathcal{A}$ models the relationship between the stress and the modified strain.
To analyze the linear stress gradient elasticity equation, we introduce the variational formulation of (1.1) in a fixed form. Mathematically speaking, the variational formulation of the linear stress gradient model VariationLSG can be regarded as a perturbed Hellinger-Reissner elasticity variation. Let $\mathbb{S}^{2}$ be the space of two-dimensional symmetric matrices, $\mathbb{R}^{2}$ be the space of two-dimensional vectors. Given $\Sigma\subset H^{1}(\Omega;\mathbb{S}^{2})$ and vector-valued space $Q\subset L^{2}(\Omega;\mathbb{R}^{2})$, the variational framework seeks solution $\sigma\in\Sigma,~u\in Q$ such that
(1.2) $\begin{cases}&\iota^{2}(\nabla\mathcal{A}\sigma,\nabla\tau)+(\mathcal{A}\sigma,\tau)+(\mathrm{div}\tau,u)=0,~\forall\tau\in\Sigma,\\ &(\mathrm{div}\sigma,v)=(f,v),~\forall v\in Q.\end{cases}$
For simplicity, here we only consider the planar case with the natural boundary conditions, i.e., $\Gamma_{f}=\emptyset.$ In this case, we set $\Sigma:=H^{1}(\Omega;\mathbb{S}^{2})$ and $Q:=L^{2}(\Omega;\mathbb{R}^{2})$. The stress and displacement can be put in to the last two slots of the smoothed stress complex:
(1.3) $\mathbf{P}_{1}~\hookrightarrow~H^{3}(\Omega)\xrightarrow{\operatorname{curl}\operatorname{curl}^{T}}H^{1}(\Omega;\mathbb{S}^{2})\xrightarrow{\operatorname{div}}L^{2}(\Omega;\mathbb{R}^{2})\to 0,$
where $\operatorname{curl}\operatorname{curl}^{T}$ denotes the planar Airy tensor. The two parameters $\lambda$ and $\iota$ are essential in modeling. When $\lambda\to+\infty$, Poisson’s ratio $\nu=\frac{\lambda}{2(\mu+\lambda)}\to\frac{1}{2}$, the materials become incompressible. When $\iota\to 0$, the effects introduced by the gradient are eliminated, and the model will converge to the standard linear elasticity model. Therefore, the goal of this paper is to provide $\lambda$- and $\iota$-robust numerical schemes.
It should be highlighted that the stress gradient elasticity is selected as one of the model problems in gradient elasticity theory. In fact, gradient elasticity theory encompasses various formulations. Notably, the strain gradient model adopts a complementary constitutive law, expressed as $\mathcal{E}-\iota^{2}\Delta\mathcal{E}=\mathcal{A}\sigma$ LSG1992 . This formulation results in a variational framework that yields a singularly perturbed fourth-order equation in the primal form Aifantis1999 ; Ming2017 ; MingH2korn . The numerical treatment can be found in MingMixLSG ; HuangLSGSIAM ; Huang2025LSG . In contrast, the stress gradient model considered in this paper produces a perturbed equation in mixed form, which is not fully addressed in the literature. Therefore, the analysis and the numerical scheme of the linear stress gradient elasticity problem can also contributes to the understanding of singular perturbation mixed problems.
Contributions. This paper develops stable finite element pairs for the stress gradient elasticity model (1.2), ensuring parameter-robust error estimates with respect to $\lambda$ and $\iota$. Especially, the results are well-adapted when $\iota=0$, leading to some stable pairs for the linear elasticity with $H^{1}$ conforming stress.
By the Ladyzhenskaya–Babuška–Brezzi conditions, it then suffices to consider $\Sigma_{h}\times Q_{h}\subset H^{1}(\Omega;\mathbb{S}^{2})\times L^{2}(\Omega;\mathbb{R}^{2})$ such that (1) (closedness) $\operatorname{div}\Sigma_{h}=Q_{h}$, and for any $q_{h}\in Q_{h}$, (2) (existence of Fortin operator) there exists $\sigma_{h}\in\Sigma_{h}$ such that $\operatorname{div}\sigma_{h}=q_{h}$ and $\|\sigma_{h}\|_{H^{1}}\lesssim\|q_{h}\|_{L^{2}}$. Throughout this paper, we call the pair is stable if the above conditions hold. Generally speaking, the goal can be separated into two parts:
  1. (1)
    When $\operatorname{div}:\Sigma_{h}\to Q_{h}$ is surjective. That is, $\operatorname{div}\Sigma_{h}=Q_{h}$.
  2. (2)
    When can we find the bounded inverse for the pair $(\Sigma_{h},\operatorname{div}\Sigma_{h}$).
Let us first introduce the candidate spaces for discretization. The most natural choice is to use continuous piecewise $P_{k}$ spaces $\Sigma_{k}^{0}$ (referred as Continuous Galerkin elements or Lagrange elements) as the discrete stress spaces, and to use discontinuous piecewise $P_{k-1}$ spaces $Q_{k-1}^{-1}$ (referred as Discontinuous Galerkin elements) as the discrete displacement spaces, to meet the regularity requirement. Moreover, the CG-DG pairs also yield great flexibility in implementation, leading to computational efficiency. The starting point of our investigation is then to design a stable CG-DG pair.
This paper establishes the first results for the stable pair $H^{1}(\mathbb{S}^{2})\times L^{2}(\mathbb{R}^{2})$ based on the CG-DG pair, providing a foundation for designing stable pairs. Specifically, we prove that if for any interior vertex, all of its edges lie in not less than four lines, then $\operatorname{div}:\Sigma_{k}^{0}\to Q_{k-1}^{-1}$ is surjective and a bounded right inverse exists, provided $k\geq 7$, see Theorem 3.1 for the precise statement. The violation can happen in two scenarios, one is the criss-cross, and the other is three-line-cross, see Figure 1 for an illustration. The presence of two types of singularities causes rank deficiencies in the divergence operator, affecting stability of the finite element pairs. Notably, the type II singularities seems less discussed to the mixed finite element methods.
Figure 1. Type I (criss-cross) and Type II singularities (three-line-cross).
In fact, the space $\operatorname{div}\Sigma_{k}^{0}$ can be explicitly characterized. In the absence of nearly singular vertices (defined precisely later), $(\Sigma_{k}^{0},\operatorname{div}\Sigma_{k}^{0})$ constitutes a stable pair, see Theorem 3.4. In practice, several approaches can generate triangular meshes avoiding singular vertices. For example, a general triangular mesh can be refined using the Morgan-Scott split, while a general rectangular mesh can be transformed into a fish-bone mesh, as illustrated in Figure 2.
Figure 2. Morgan–Scott split and fish-bone mesh.
On the other hand, we are interested in finding the finite element pair satisfying the unconditional stability. Here “unconditional” indicates that the inf-sup condition can be established without any specific geometric structures on triangulations (in this case, singular vertices), and only depends on the shape regularity. Motivated by the Falk–Neilan Stokes pair FalkNeilan13 , CG-DG pairs with extra vertex continuity are considered. In Theorem 3.2, we show that the pair $\Sigma_{k}^{2}\times Q_{k-1}^{1}$ achieves unconditional stability. Here $\Sigma_{k}^{2}:=\{\sigma\in\Sigma_{k}^{0},\sigma\text{ is }C^{2}\text{ at vertices}\}$, and $Q_{k-1}^{1}:=\{\sigma\in Q_{k}^{-1},\sigma\text{ is }C^{1}\text{ at vertices}\}$. The extra smoothness at vertices cannot be reduced Theorem 3.3. The intuition of the unconditional stability comes from the fact that $C^{4}$ continuity is required for constructing $H^{3}$ conforming finite elements HuLinWuYuan25 .
Relationship to the Stokes complex. The results presented above share some similarities with respect to the divergence-free Stokes pair. Correspondingly, (1.3) has strong relationship with the following Stokes complex:
(1.4) $\mathbf{P}_{0}~\hookrightarrow~H^{2}(\Omega)\xrightarrow{\operatorname{curl}}H^{1}(\Omega;\mathbb{R}^{2})\xrightarrow{\operatorname{div}}L^{2}(\Omega)\to 0.$
However, the motivations of the study are slightly different from each other. Unlike the Stokes pair, where weak divergence-free conditions suffice, the stress gradient model requires exact divergence surjectivity due to the non-coercive bilinear form. As a consequence, the surjectivity and the complex structure yield much importance for smoothed stress complexes.
We now compare the results from the divergence-free Stokes pair and the results in this paper. From the comparison, we can see the algebraic patterns. The CG-DG pair in the Stokes equation is first constructed and analyzed by Scott and Vogelius ScottVogelius85 , and a recent re-investigation can be found in GuzmanScott19 . They showed that the CG-DG pair is stable if there is no criss-cross internal vertices and the polynomial degree $k\geq 4$.
Building on the discrete Stokes complex GuzmanNeilan14 ; FalkNeilan13 ; Neilan15 ; HuZhangZhang22 , stable pairs with divergence-free properties can be constructed. Specifically, the Hermite element (globally continuous with $C^{1}$ vertex continuity) can be chosen as the discrete velocity space, paired with a globally discontinuous space with $C^{0}$ continuity for the discrete pressure space. However, we can show that this is insufficent for (1.2). In fact, we can prove that for such a Hermite pair in (1.2), the divergence operator is onto if and only if there is no criss-cross singular vertices. A comprehensive comparison of the results are presented in Figure 3.
Stokes Stress Gradient Stokes Stress Gradient
Type I Type I Type II Type II
$(\Sigma_{k}^{0},Q_{k-1}^{-1})$ 1 3 0 1
$(\Sigma_{k}^{1},Q_{k-1}^{0})$ 0 1 0 0
$(\Sigma_{k}^{2},Q_{k-1}^{1})$ 0 0 0 0
Figure 3. The rank deficiency of Stokes pair and stress gradient pair.
Notations. In this paper, let $\mathcal{T}$ be a conforming planar triangulation of a polygon domain $\Omega$. Moreover, we assume that $\Omega$ is contractible. Let $\mathcal{E}$ and $\mathcal{V}$ be the set of all edges and vertices, respectively. We will also use $T$, $e$, $v$ to denote a face (element), edge, and vertex.
Define the rigid body motion space $\mathbf{RM}=\mathrm{span}\{(1,0)^{T},~(0,1)^{T},~(-y,x)^{T}\}$, which is the kernel of the deformation tensor. We define the perpendicular operator for a vector $\begin{bmatrix}a\\ b\end{bmatrix}$ as $\begin{bmatrix}a\\ b\end{bmatrix}^{\perp}:=\begin{bmatrix}b\\ -a\end{bmatrix}$.
Organization. The paper is organized as follows. In Section 2, we discuss the well-posedness and a priori estimates of (1.2). In Section 3, we show the main results of the construction of the stable pairs, and the corresponding error estimates are shown in Section 3.4. In Section 4, we show the numerical results. Some concluding remarks and the discussions are shown in Section 5.
2. Wellposedness and a priori estimates for continuous problems
In this section, we show the well-posedness of (1.2) and some parameter-robust estimates. For $\iota>0$, let $a_{\iota}(\sigma,\tau):=\iota^{2}(\nabla\mathcal{A}\sigma,\nabla\tau)+(\mathcal{A}\sigma,\tau)$ and $b(\sigma,u)=(\mathrm{div}\sigma,u).$ Therefore, the mixed form can be reformulated as
(2.1) $\begin{cases}a_{\iota}(\sigma,\tau)+b(\tau,u)=0,&\forall\tau\in H^{1}(\Omega;\mathbb{S}^{2}),\\ b(\sigma,v)=(f,v),&\forall v\in L^{2}(\Omega;\mathbb{R}^{2}).\end{cases}$
For the stress space $\Sigma:=H^{1}(\Omega;\mathbb{S}^{2})$, we define the $\iota$-norm as $\|\sigma\|_{\iota}:=\iota|\sigma|_{H^{1}}+\|\mathrm{div}\sigma\|_{L^{2}}+\|\sigma\|_{L^{2}}$. For the displacement space $Q:=L^{2}(\Omega;\mathbb{R}^{2})$, we equip it with the standard $L^{2}$ norm. The goal of this section is to derive some estimates independent of $\lambda$ and $\iota$.
To verify the wellposedness, we need to show that $a_{\iota}$ is coercive in the kernel space $Z:=\{\sigma\in H^{1}(\Omega;\mathbb{S}^{2})~:~\mathrm{div}\sigma=0\}$. We recall some results from the standard analysis of the linear elasticity equations. Hereafter, the hidden constants of $\lesssim$ are independent of $\lambda$ and $\iota$.
Lemma 2.1.
For $\tau\in H^{1}(\Omega;\mathbb{S}^{2})$, the estimates hold pointwise:
(2.2) $\nabla\tau^{D}:\nabla\tau^{D}+\mathrm{div}\tau\cdot\mathrm{div}\tau\geq\frac{1}{8}\nabla\tau:\nabla\tau.$
Proof.
Denote by $\|v\|_{0}=v\cdot v$ and $|v|_{1}=\nabla v:\nabla v$ as the pointwise $L^{2}$ norm and $H^{1}$ semi-norm. By direct computation, we have
(2.3) $\nabla\tau^{D}:\nabla\tau^{D}=\frac{1}{2}|\tau_{11}-\tau_{22}|_{1}^{2}+2|\tau_{12}|_{1}^{2}=\frac{1}{2}|\tau_{11}|_{1}^{2}+\frac{1}{2}|\tau_{22}|_{1}^{2}+2|\tau_{12}|_{1}^{2}+\nabla\tau_{11}\cdot\nabla\tau_{22}.$
Since we have
$\tau_{11,x}\tau_{22,x}+\tau_{11,y}\tau_{22,y}=(\tau_{11,x}+\tau_{12,y})\tau_{22,x}+\tau_{11,y}(\tau_{12,x}+\tau_{22,y})-\tau_{12,y}\tau_{22,x}-\tau_{12,x}\tau_{11,y}.$
Therefore,
$\nabla\tau_{11}\cdot\nabla\tau_{22}\geq-\|\operatorname{div}\tau\|_{0}^{2}-\frac{1}{8}|\tau_{22}|_{1}^{2}-\frac{1}{8}|\tau_{11}|_{1}^{2}-2|\tau_{12}|_{1}^{2}-\frac{1}{4}|\tau_{11}|_{1}^{2}-\frac{1}{4}|\tau_{22}|_{1}^{2}.$
Therefore, we finish the proof.
A corollary of Equation 2.2 is
$\|\nabla\tau\|^{2}_{L^{2}}\leq 8\|\nabla\tau^{D}\|^{2}_{L^{2}}+8\|\operatorname{div}\tau\|^{2}_{L^{2}}.$
We now introduce the following estimates for symmetric matrix-valued function $\tau$:
(2.4) $\mathcal{A}\tau:\tau=(\frac{1}{2\mu}\tau^{D}:\tau^{D}+\frac{1}{4\lambda+4\mu}\mathrm{tr}(\tau)\mathrm{tr}(\tau)),~\forall\tau\in\mathbb{S}^{2},$
where $\tau^{D}=\tau-\frac{1}{2}\mathrm{tr}\tau\mathrm{I}$ is the traceless part of $\tau$.
Proposition 2.1 (Coercivity).
For all $\tau\in Z$ (which implies that $\operatorname{div}\tau=0$) and $\int_{\Omega}\mathrm{tr}(\tau)=0,$
(2.5) $\iota^{2}(\nabla\mathcal{A}\tau,\nabla\tau)+(\mathcal{A}\tau,\tau)\gtrsim\|\tau\|^{2}_{\iota}.$
Proof.
The following result can be found in (MixedFEMbook, , Proposition 9.1.1): For $\tau\in L^{2}(\Omega;\mathbb{S}^{2})$ and $\int_{\Omega}\mathrm{tr}(\tau)=0$, it holds that
$\|\tau\|_{L^{2}}\lesssim(\|\tau^{D}\|_{L^{2}}+\|\operatorname{div}\tau\|_{L^{2}}).$
Here the hidden constant only depends on $\Omega$. Together with (2.4), we have
$\|\tau\|_{L^{2}}\lesssim\|\tau^{D}\|_{L^{2}}\leq 2\mu(\mathcal{A}\tau,\tau),$
when $\tau\in Z$ and $\int_{\Omega}\mathrm{tr}(\tau)=0$. Applying gradient operator to (2.4), we obtain
$\|\frac{1}{8}\nabla\tau\|_{L^{2}}\leq\|\nabla\tau^{D}\|_{L^{2}}\leq 2\mu(\nabla\mathcal{A}\tau,\nabla\tau),~\forall\tau\in Z.$
Thus we finish the proof. ∎
The following lemma gives the inf-sup condition of $b(\sigma,v)$, which is a standard result for linear elasticity equations MixedFEMbook .
Lemma 2.2 (Inf-sup conditions).
For any $v\in L^{2}(\Omega;\mathbb{R}^{2})$, there exists a $\tau\in H^{1}(\Omega;\mathbb{S}^{2})$ such that
$\mathrm{div}\tau=v,~\|\tau\|_{\iota}\lesssim\|v\|_{L^{2}}.$
Proof.
For all $v\in L^{2}(\Omega;\mathbb{S}^{2})$, there exists a exact solution $u\in H^{2}(\Omega)\cap H_{0}^{1}(\Omega)$, such that
$-\operatorname{div}(2\mu\operatorname{sym}\operatorname{grad}(u)+\lambda\operatorname{div}uI)=v,$
and $\|u\|_{H^{2}}+\lambda\|\operatorname{div}u\|_{L^{2}}\leq\|v\|_{L^{2}}$. Let
$\tau=2\mu\operatorname{sym}\operatorname{grad}(u)+\lambda\operatorname{div}uI\in\Sigma.$
Since $\|\tau\|_{\iota}\leq 2\|\tau\|_{H^{1}}$ we finish the proof. ∎
By the Ladyzhenskaya-Babuška-Brezzi theory, we establish the wellposedness of problem (1.2).
Theorem 2.1.
For all $f\in L^{2}(\Omega;\mathbb{R}^{2})$, there exists a unique solution pair $(\sigma_{\iota},u_{\iota})\in H^{1}(\Omega;\mathbb{S}^{2})\times L^{2}(\Omega;\mathbb{R}^{2})$ of (1.2), such that
$\|\sigma_{\iota}\|_{\iota}+\|u_{\iota}\|_{L^{2}}\lesssim\|f\|_{L^{2}}.$
Here the hidden constant is independent of $\iota$ and $\lambda$.
Proof.
Since $\|\sigma\|_{\iota}\geq\|\operatorname{div}\sigma\|_{L^{2}}$, the boundedness property of $b(\sigma,v)$ is trivial. By inf-sup stability, there exists unique $\sigma^{\perp}\in Z^{\perp}$, the orthogonal complement of $Z$ in $H^{1}(\Omega;\mathbb{S}^{2})$ with respect to $\iota$ norm, such that
$b(\sigma^{\perp},v)=(f,v),~\forall v\in Q,\text{ and }\|\sigma^{\perp}\|_{\iota}\lesssim\|f\|_{L^{2}}.$
By (2.4), and $\|\mathrm{tr}(\sigma)\|_{L^{2}}\leq 2\|\sigma\|_{L^{2}},$ we have
$a_{\iota}(\sigma,\tau)\leq(\frac{1}{2\mu}+\frac{1}{(\lambda+\mu)})(\|\sigma\|_{L^{2}}\|\tau\|_{L^{2}}+\iota^{2}\|\nabla\sigma\|_{L^{2}}\|\nabla\tau\|_{L^{2}}),~\forall~\tau,\sigma\in\Sigma.$
By Lax-Milgram Theorem, there exists $\sigma^{0}\in Z$, such that
$a_{\iota}(\sigma^{0},\tau_{z})=-a_{\iota}(\sigma^{\perp},\tau_{z}),~\forall~\tau_{z}\in Z,$
and $\|\sigma^{0}\|_{\iota}\lesssim\|\sigma^{\perp}\|_{\iota}.$ Then $\sigma_{\iota}=\sigma^{\perp}+\sigma^{0}$ is the unique solution of (1.2). Finally, by the inf-sup condition again and the first equation of (1.2), there exists $u_{\iota}\in L^{2}(\Omega;\mathbb{R}^{2})$, such that
$b(\tau,u_{\iota})=-a_{\iota}(\sigma_{\iota},\tau),~\forall~\tau\in Z^{\perp},$
and $\|u_{\iota}\|_{L^{2}}\lesssim\|\sigma_{\iota}\|_{\iota}.$ Thus we finish the proof. ∎
Now we derive the asymptotic result of the solution. Let $(\sigma_{\iota},u_{\iota})$ solves (1.2), and $(u_{0},\sigma_{0})$ solves the following equation:
(2.6) $\begin{cases}(\mathcal{A}\sigma_{0},\tau)+b(\tau,u_{0})=0,&\forall\tau\in H(\operatorname{div};\Omega),\\ b(\sigma_{0},v)=(f,v),&\forall v\in L^{2}(\Omega).\end{cases}$
Proposition 2.2.
Assume $\sigma_{0}\in H^{2}(\Omega;\mathbb{S}^{2}),~u_{0}\in H^{2}(\Omega;\mathbb{R}^{2}),$ then we have $\|\sigma_{\iota}-\sigma_{0}\|_{\iota}+\|u_{\iota}-u_{0}\|_{L^{2}}=O(\iota^{3/2}).$
Proof.
Clearly, for all $\tau\in H^{1}(\Omega;\mathbb{S}^{2})$ and $v\in L^{2}(\Omega;\mathbb{R}^{2})$ we have
(2.7) $\begin{cases}\iota^{2}(\nabla\mathcal{A}(\sigma_{\iota}-\sigma_{0}),\nabla\tau)+(\mathcal{A}(\sigma_{\iota}-\sigma_{0}),\tau)+b(\tau,u_{\iota}-u_{0})=-\iota^{2}(\nabla\mathcal{A}\sigma_{0},\nabla\tau)\\ b(\sigma_{\iota}-\sigma_{0},v)=0.\end{cases}$
Thus, we have $\operatorname{div}(\sigma_{\iota}-\sigma_{0})=0.$ Let $\tau=\sigma_{\iota}-\sigma_{0},$ we then have
$\iota^{2}(\nabla\mathcal{A}(\sigma_{\iota}-\sigma_{0}),\nabla(\sigma_{\iota}-\sigma_{0}))+(\mathcal{A}(\sigma_{\iota}-\sigma_{0}),\sigma_{\iota}-\sigma_{0})=-\iota^{2}(\nabla\mathcal{A}\sigma_{0},\nabla(\sigma_{\iota}-\sigma_{0})).$
This implies
(2.8) $\begin{split}\iota^{2}|\sigma_{\iota}-\sigma_{0}|_{H^{1}}^{2}+\|\sigma_{\iota}-\sigma_{0}\|_{L^{2}}^{2}\lesssim&-\iota^{2}(\nabla\mathcal{A}\sigma_{0},\nabla(\sigma_{\iota}-\sigma_{0}))\\ =&\iota^{2}(\Delta\mathcal{A}\sigma_{0},\sigma_{\iota}-\sigma_{0})-\iota^{2}\int_{\partial\Omega}\partial_{n}(\mathcal{A}\sigma_{0}):(\sigma_{\iota}-\sigma_{0})\\ \lesssim&\iota^{2}\|\partial_{n}\sigma_{0}\|_{L^{2}(\partial\Omega)}\|\sigma_{\iota}-\sigma_{0}\|_{L^{2}(\partial\Omega)}+\iota^{2}|\sigma_{0}|_{H^{2}}\|\sigma_{\iota}-\sigma_{0}\|_{L^{2}}.\end{split}$
By trace theorem, we then have
$\iota^{2}|\sigma_{\iota}-\sigma_{0}|_{H^{1}}^{2}+\|\sigma_{\iota}-\sigma_{0}\|_{L^{2}}^{2}\lesssim\iota^{2}\|\partial_{n}\sigma_{0}\|_{L^{2}(\partial\Omega)}\|\sigma_{\iota}-\sigma_{0}\|_{H^{1}}^{1/2}\|\sigma_{\iota}-\sigma_{0}\|_{L^{2}}^{1/2}+\iota^{2}|\sigma_{0}|_{H^{2}}\|\sigma_{\iota}-\sigma_{0}\|_{L^{2}}.$
By Cauchy-Schwarz inequality, we then have
$\iota^{2}|\sigma_{\iota}-\sigma_{0}|_{H^{1}}^{2}+\|\sigma_{\iota}-\sigma_{0}\|_{L^{2}}^{2}\lesssim\iota^{3}\|\partial_{n}\sigma_{0}\|_{L^{2}(\partial\Omega)}^{2}+\iota^{4}|\sigma_{0}|_{H^{2}}^{2},$
which is equivalent to
$\iota|\sigma_{\iota}-\sigma_{0}|_{H^{1}}+\|\sigma_{\iota}-\sigma_{0}\|_{L^{2}}\lesssim\iota^{3/2}\|\partial_{n}\sigma_{0}\|_{L^{2}(\partial\Omega)}+\iota^{2}|\sigma_{0}|_{H^{2}}.$
Next, we give the estimates on the term $u_{\iota}-u_{0}.$ Take $\xi\in H^{1}(\Omega;\mathbb{S}^{2})$ such that $\|\xi\|_{H^{1}}\lesssim\|u_{\iota}-u\|_{L^{2}}$ and $\operatorname{div}\xi=u_{\iota}-u_{0}$. Therefore, we have
$\iota^{2}(\nabla\mathcal{A}(\sigma_{\iota}-\sigma_{0}),\nabla\xi)+(\mathcal{A}(\sigma_{\iota}-\sigma_{0}),\xi)+\|u_{\iota}-u_{0}\|_{L^{2}}=-\iota^{2}(\nabla\mathcal{A}\sigma_{0},\nabla\xi).$
By Cauchy-Schwarz inequality, we then have
$\|u_{\iota}-u_{0}\|_{L^{2}}^{2}\lesssim\iota^{2}|\sigma_{\iota}|_{H^{1}}\|\xi\|_{H^{1}}+\|\sigma_{\iota}-\sigma_{0}\|_{L^{2}}\|\xi\|_{L^{2}},$
which implies
$\|u_{\iota}-u_{0}\|_{L^{2}}\lesssim\iota^{2}|\sigma_{\iota}|_{H^{1}}+\|\sigma_{\iota}-\sigma_{0}\|_{L^{2}}\lesssim\iota^{3/2}\|\partial_{n}\sigma_{0}\|_{L^{2}{\partial\Omega}}+\iota^{2}|\sigma_{0}|_{H^{1}}+\iota^{3}|\sigma_{0}|_{H^{2}}.$
3. Stable Finite Element Pairs for stress gradient elasticity
This section constructs a family of stable pairs for stress gradient elasticity. Specifically, we seek a stress-displacement pair $(\Sigma_{h},Q_{h})\subseteq H^{1}(\Omega;\mathbb{S}^{2})\times L^{2}(\Omega;\mathbb{R}^{2})$ such that the divergence operator $\operatorname{div}:\Sigma_{h}\to Q_{h}$ is surjective and admits a bounded (generalized) inverse. That is, for each $q_{h}\in Q_{h}$, there exists $\sigma_{h}\in\Sigma_{h}$ satisfying $\operatorname{div}\sigma_{h}=q_{h}$ and $\|\sigma_{h}\|_{H^{1}}\lesssim\|q_{h}\|_{L^{2}}$. For such a pair, we can consider the discrete problem
(3.1) $\begin{cases}a_{\iota}(\sigma_{h},\tau_{h})+b(\tau_{h},u_{h})=0,&\forall\tau_{h}\in\Sigma_{h},\\ b(\sigma_{h},v_{h})=(f,v_{h}),&\forall v_{h}\in Q_{h}.\end{cases}$
The wellposedness and the error estimates of the discrete problem are discussed in Section 3.4.
In this section, we introduce the CG-DG pairs as follows. Define the following stress spaces
$\Sigma^{0}_{k}=\{\sigma\in C^{0}(\mathcal{T};\mathbb{S}^{2}):\sigma|_{T}\in P_{k}(T;\mathbb{S}^{2}),\sigma\in C^{0}(\mathcal{V})\},$
$\Sigma^{1}_{k}=\{\sigma\in C^{0}(\mathcal{T};\mathbb{S}^{2}):\sigma|_{T}\in P_{k}(T;\mathbb{S}^{2}),\sigma\in C^{1}(\mathcal{V})\},$
$\Sigma^{2}_{k}=\{\sigma\in C^{0}(\mathcal{T};\mathbb{S}^{2}):\sigma|_{T}\in P_{k}(T;\mathbb{S}^{2}),\sigma\in C^{2}(\mathcal{V})\}.$
Here, $\sigma\in C^{0}(\mathcal{V})$ means that $\sigma$ is continuous at vertices, $\sigma\in C^{r}(\mathcal{V})$ means that $\nabla^{r^{\prime}}\sigma$ is continuous at vertices for $0\leq r^{\prime}\leq r.$
Accordingly, we define the following displacement space.
$Q^{-1}_{k-1}=\{q\in L^{2}(\mathcal{T};\mathbb{R}^{2}):q|_{T}\in P_{k-1}(T;\mathbb{R}^{2})\},$
$Q^{0}_{k-1}=\{q\in L^{2}(\mathcal{T};\mathbb{R}^{2}):q|_{T}\in P_{k-1}(T;\mathbb{R}^{2}),q\in C^{0}(\mathcal{V})\},$
$Q^{1}_{k-1}=\{q\in L^{2}(\mathcal{T};\mathbb{R}^{2}):q|_{T}\in P_{k-1}(T;\mathbb{R}^{2}),q\in C^{1}(\mathcal{V})\}.$
To address the rank deficiency and the stability issue caused by the geometric structure, we now formally define two types of singularities.
Definition 3.1 (Type I Singularity).
A vertex $v$ is said to have a Type I singularity if and only if it is an interior vertex and all edges incident to $v$ lie on at most two distinct lines.
Such singularities have been thoroughly investigated in the Stokes’ pair ScottVogelius85 ; GuzmanScott19 , and the construction of the $C^{1}$ spline space on general triangular meshes MorganScott75 .
The second type of singularity is the three-line-cross singularity (Type II), which is not fully addressed in the literature.
Definition 3.2 (Type II Singularity).
A vertex $v$ is said to have a Type II singularity if and only if it is an interior vertex and all edges incident to $v$ lie on exactly three distinct lines. A singular vertex is called non-degenerate if it is surrounded by six elements (see Figure 4); otherwise, it is called degenerate (see Figure 5).
Figure 4. Non-degenerate Type II singularity.
Figure 5. Degenerate Type II singularity. The numbers of the surrounding element of a Type II singular vertex can be 5,4,3. We denote them as Type II-1, II-2, and II-3, accordingly.
In this section, we establish a key result on the surjectivity of the divergence operator for the stress-displacement pairs proposed above in stress gradient elasticity. The main theorem presented below characterizes the precise geometric conditions under which the divergence operator $\operatorname{div}:\Sigma_{h}\to Q_{h}$ is surjective, ensuring the existence of a generalized inverse operator.
Theorem 3.1.
Suppose that $k\geq 7$. Then the following statements hold:
  1. (1)
    (Unconditional Stability) For the pair $(\Sigma_{k}^{2},Q_{k-1}^{1})$, the divergence operator $\operatorname{div}:\Sigma_{k}^{2}\to Q_{k-1}^{1}$ is surjective for any triangulation $\mathcal{T}$.
  2. (2)
    For the Hermite pair $(\Sigma_{k}^{1},Q_{k-1}^{0})$, the divergence operator $\operatorname{div}:\Sigma_{k}^{1}\to Q_{k-1}^{0}$ is surjective for any triangulation $\mathcal{T}$ that has no Type I singularities.
  3. (3)
    For the Lagrange pair $(\Sigma_{k}^{0},Q_{k-1}^{-1})$, the divergence operator $\operatorname{div}:\Sigma_{k}^{0}\to Q_{k-1}^{-1}$ is surjective for any triangulation $\mathcal{T}$ that has no Type I or Type II singularities.
For a general triangulation $\mathcal{T}$, the following quantitative results hold:
  1. (2’)
    The rank deficiency of the pair $(\Sigma_{k}^{1},Q_{k-1}^{0})$ is given by
    (3.2) $\dim Q_{k-1}^{0}-\dim\operatorname{div}\Sigma_{k}^{1}=|\mathcal{V}_{I}|.$
  2. (3’)
    The rank deficiency of the pair $(\Sigma_{k}^{0},Q_{k-1}^{-1})$ is given by
    (3.3) $\dim Q_{k-1}^{-1}-\dim\operatorname{div}\Sigma_{k}^{0}=|\mathcal{V}_{II}|+3|\mathcal{V}_{I}|.$
Here, $\mathcal{V}_{I}$ denotes the set of vertices with Type I singularities, and $\mathcal{V}_{II}$ denotes the set of vertices with Type II singularities. See Figure 3 for a summary and comparison with the Stokes pair.
Proof.
(1) comes from Theorem 3.2. (2) and (2’) come from Theorem 3.3. (3) and (3’) come from Theorem 3.4. ∎
Remark 3.1.
The polynomial degree requirement $k\geq 7$ is imposed for technical reasons. In this paper, our primary focus is to classify singular vertices, while determining the optimal polynomial degree requirement is not our main objective at present and will be addressed in future work. Especially, reducing the degrees of freedom may require additional conditions, see GuzmanScott18 .
Remark 3.2.
We make a comparison of the proposed pairs with the Stokes pairs. The pairs of (2) and (2’) can be regarded as matrix-valued versions of the Falk–Neilan Stokes pair, while the results from items (3) and (3’) can be regarded as matrix-valued versions of the Scott–Vogelius Stokes pair. This can help the readers compare the results better.
Sketch of Proofs
The proof of the theorem above relies on a detailed characterization of $\operatorname{div}\Sigma_{k}^{r}$. In the following, we use three subsections to provide this characterization and establish the stability results for the pair $(\Sigma_{k}^{r},\operatorname{div}\Sigma_{k}^{r})$ for $r=2,1,0$, respectively.
The theorem will be obtained item by item, and for clarity we show the sketch of the overall proof here. To establish unconditional stability, we construct the discrete smoothed stress complexes, starting from $C^{2}$ scalar element spaces. Based on the bubble complexes and the finite element complexes, we can derive the unconditional stability of the global function space pair $(\Sigma_{k}^{2},Q_{k}^{1})$. To remove the extra continuity at vertices, we consider the function data on the local vertex patches, and make suitable modifications to proceed. The geometric constraints appear during the modification: When dealing with $\Sigma_{k}^{1}$, only type I singularities appear; when dealing with $\Sigma_{k}^{0}$, both type I and type II singularities appear.
Therefore, the proof relies on analyzing data on the vertex patches, calling for more notations. For a vertex $z\in\mathcal{V}$, denote by $\operatorname{star}(z)$ the patch of $z$, namely, it consists of elements surrounding $z$. The elements $T_{i}$ in $\operatorname{star}(z)$ are numbered counterclockwise as $T_{1},T_{2},\dots,T_{m}$, where $m$ is the total number of elements in the patch. We use $y_{s}$ to denote the vertex shared by elements $T_{s}$ and $T_{s+1}$, and $t_{s}$ ($n_{s}:=t_{s}^{\perp}$, resp.) to represent the unit tangent (normal, resp.) vector of the directed edge $\overrightarrow{zy_{s}}$. When the vertex is interior, the subscription is periodically labeled. See Figure 6 for an illustration. Additionally, we define $h_{s}$ as the height with respect to vertex $y_{s}$. Specifically, we have
(3.4) $h_{s}:=\begin{cases}|\overline{zy_{s}}|\sin\theta_{s}&\text{ in }T_{s},\\ |\overline{zy_{s}}|\sin\theta_{s+1}&\text{ in }T_{s+1}.\end{cases}$
When focusing on a single element $T_{s}$, we may denote $t_{-}$ as $t_{s}$ and $t_{+}$ as $t_{s+1}$ if the context permits.
For each vertex $z\in\mathcal{V}$, we define the nodal basis function as
$\psi_{z}(v)=\begin{cases}1&\text{if }v=z,\\ 0&\text{if }v\neq z.\end{cases}$
Within each element $T$, the restriction of $\psi_{z}$ to $T$ is the barycentric coordinate function associated with vertex $z$. Similarly, we use $n_{\pm}$ and $\psi_{\pm}$ when the context permits.
[…]
Figure 6. An illustration for the notations on the patch $z$.
Using the notations introduced above, we can define a measure to quantify the extent to which a triangulation is singular or nearly singular with respect to Type I or Type II singularities. For any interior vertex $x\in\mathcal{V}$, we define
(3.5) $\Theta_{I}(x):=\max\{|\sin(\theta_{i}+\theta_{i+1})|\},$
where the maximum is taken over all $i$ in the patch $\operatorname{star}x$. It can be verified that $\Theta_{I}(x)=0$ if and only if $x\in\mathcal{V}_{I}$. Similarly, we define
(3.6) $\Theta_{II}(x):=\max\{\min(|\sin(\theta_{i}+\theta_{i+1}+\theta_{i+2})|,|\sin(\theta_{i}+\theta_{i+1})|,|\sin(\theta_{i+1}+\theta_{i+2})|)\},$
for any interior vertex $x\in\mathcal{V}$, where the maximum is taken over all $i$ in the patch $\operatorname{star}x$. It can be verified that $\Theta_{II}(x)=0$ if and only if $x\in\mathcal{V}_{I}$ or $x\in\mathcal{V}_{II}$.
In this section, the hidden constants in $\lesssim$ might depend on the region $\Omega$, the polynomial degree $k$, and the shape regularity constant, but independent with the triangulation $\mathcal{T}$. The dependency on singularities in triangulations will be given explicitly via $\Theta_{I}(x)$ or $\Theta_{II}(x)$.
Remark 3.3.
The quantities mentioned above are defined only for the interior vertices. For the boundary vertices, they do not contribute to the cokernels of divergence operators.
3.1. The Pair $\Sigma_{k}^{2}\times Q_{k-1}^{1}$: Unconditional Stability and Discrete Complexes
In this subsection, we show that the pair $\Sigma_{k}^{2}\times Q_{k-1}^{1}$ is unconditionally stable with respect to both types of singularities. The main theorem is summarized as follows.
Theorem 3.2.
The divergence operator $\operatorname{div}:\Sigma_{k}^{2}\to Q_{k-1}^{1}$ is surjective for any mesh $\mathcal{T}$, provided that $k\geq 7$. Moreover, if $\mathcal{T}$ is shape regular, then for each $q\in Q_{k-1}^{1}$, there exists $\sigma\in\Sigma_{k}^{2}$ such that $\operatorname{div}\sigma=q$ and $\|\sigma\|_{H^{1}}\lesssim\|q\|_{L^{2}}.$
Theorem 3.2 establishes that unconditional stability is achieved by imposing $C^{2}$ continuity at the vertices in the stress element space $\Sigma_{k}^{2}$ and $C^{1}$ continuity at the vertices in the displacement element space $Q_{k-1}^{1}$.
This theorem arises from a finite element discretization of the smoothed stress complex (1.3). To elucidate this, we recall the $H^{3}$ conforming finite element $U_{k+2}^{4}$ from Zenisek74 ; HuLinWu24 , defined for $k\geq 7$. The local shape function space of $U_{k+2}^{4}$ is the polynomial space $P_{k+2}$. For $u\in P_{k+2}$ on an element $T$, the degrees of freedom are given as follows:
  1. (1)
    $\nabla^{\alpha}u(v)$ for each vertex $v$ with $|\alpha|\leq 4$;
  2. (2)
[…]
  3. (3)
    $\int_{T}u\,p$ for $p\in P_{k-7}(T)$.
It can be readily checked that the finite element $U_{k+2}^{4}$ is unisolvent, and the resulting space is $H^{3}$ conforming.
The local shape function space of $\Sigma_{k}^{2}$ is $P_{k}(T;\mathbb{S}^{2})$, and for $\sigma\in P_{k}(T;\mathbb{S}^{2})$, the degrees of freedom are given as follows:
  1. (1)
    $\nabla^{\alpha}\sigma(v)$ for each vertex $v$ of $T$ and $|\alpha|\leq 2$;
  2. (2)
    $\int_{e}\sigma:p$ for $p\in P_{k-6}(e;\mathbb{S}^{2})$ for each edge $e$ of $T$;
  3. (3)
    $\int_{T}\sigma:p$ for $p\in P_{k-3}^{(0)}(T;\mathbb{S}^{2}).$ Here $P_{k-3}^{(0)}(T):=\{p\in P_{k-3}(T):p(v)=0,\,\,\forall v\in\mathcal{V}(T)\}.$
Then, we can construct the following sequence:
(3.7) $\mathbf{P}_{1}~{\hookrightarrow}~U^{4}_{k+2}(\mathcal{T})\xrightarrow{\operatorname{curl}\operatorname{curl}^{T}}\Sigma_{k}^{2}(\mathcal{T})\xrightarrow{\operatorname{div}}Q_{k-1}^{1}(\mathcal{T})\to 0.$
The exactness of (3.7) can be checked from direct calculation.
Proposition 3.1.
The finite element sequence (3.7) forms a complex. When $\Omega$ is topologically trivial, this sequence is exact.
Proof.
For $\sigma\in\Sigma_{k}^{2}(\mathcal{T})\subseteq H^{1}(\Omega;\mathbb{S}^{2})$ with $\operatorname{div}\sigma=0$, there exists $\psi\in H^{3}$ such that $\psi$ is piecewise polynomial. Since $\psi\in H^{3}$, we know that $\psi$ and $\nabla\psi$ are continuous at vertices. Therefore, $\psi\in U_{k+2}^{4}(\mathcal{T})$. To establish exactness, it suffices to verify the dimension count for the sequence (3.7). By direct calculation, we obtain the following dimensions:
$\dim U_{k+2}^{4}(\mathcal{T})$ $=\frac{1}{2}(k-6)(k-5)|\mathcal{T}|+(3k-18)|\mathcal{E}|+15|\mathcal{V}|,$
$\dim\Sigma_{k}^{2}(\mathcal{T})$ $=\left(\frac{3}{2}(k+2)(k+1)-9(k+1)\right)|\mathcal{T}|+(3k-15)|\mathcal{E}|+18|\mathcal{V}|,$
$\dim Q_{k-1}^{1}(\mathcal{T})$ $=(k(k+1)-18)|\mathcal{T}|+6|\mathcal{V}|.$
Since
$\dim U_{k+2}^{4}(\mathcal{T})-\dim\Sigma_{k}^{2}(\mathcal{T})+\dim Q_{k-1}^{1}(\mathcal{T})=3(|\mathcal{T}|-|\mathcal{E}|+|\mathcal{V}|)=3,$
where the final equality holds for a topologically trivial domain $\Omega$ by the Euler equation, it then follows that the finite element sequence is exact. ∎
[…]
$\mathring{\Sigma}_{k}(T)$ :=$:=\left\{\sigma\in P_{k}(T;\mathbb{S}^{2}):\nabla^{\alpha}\sigma(v)=0,\,\forall|\alpha|\leq 2,\,\sigma|_{\partial T}=0\right\},$
$\mathring{Q}_{k-1}(T)$ :=$:=\left\{q\in P_{k-1}(T;\mathbb{R}^{2}):\nabla^{\alpha}q(v)=0,\,\forall|\alpha|\leq 1\right\}.$
Proposition 3.2.
The following sequence on each element $T\in\mathcal{T}$:
[…]
forms a complex and is exact. Here $\cap\mathbf{RM}^{\perp}$ means that the $\mathbf{RM}$ moment is zero.
Proof.
It is straightforward to see (3.8) is a complex. For exactness, it suffices to verify the dimension count. We compute the dimensions of the local bubble spaces as follows:
[…]
$\dim\mathring{\Sigma}_{k}(T)$ $=3\dim P_{k}(T)-9(k-5)-54=\frac{3}{2}(k+2)(k+1)-9(k-5)-54,$
$\dim\mathring{Q}_{k-1}(T)$ $=2\dim P_{k-1}(T)-12=k(k+1)-18.$
It follows that
[…]
Therefore, we conclude the exactness. ∎
By the surjectivity result and scaling argument, we have the bubble stability.
Corollary 3.1 (Bubble Stability).
Suppose that $k\geq 7$. For any $q\in P_{k-1}(T;\mathbb{R}^{2})$ satisfying:
  1. (1)
    $q(v)=\nabla q(v)=0$ for all vertices $v\in\mathcal{V}(T)$, and
  2. (2)
[…]
there exists $\sigma\in\mathring{\Sigma}_{k}(T)$ such that $\operatorname{div}\sigma=q$ and $\|\sigma\|_{H^{1}}\lesssim\|q\|_{L^{2}},$ where the hidden constant depends only on the shape regularity constant of $T$.
Now the proof of Theorem 3.2 follows a standard argument FalkNeilan13 ; HuZhang14 , with the help of Corollary 3.1.
Proof of Theorem 3.2.
Given any $q_{h}\in Q_{k-1}^{1}\subset L^{2}(\Omega;\mathbb{R}^{2})$, it suffices to show that there exists $\tau\in H^{1}(\Omega;\mathbb{S}^{2})$ such that $\mathrm{div}\tau=q_{h}.$ The proof will be divided into two steps.
In the first step, we select $\sigma^{(1)}\in\Sigma_{k}^{2}$ such that
  1. (1)
    For all vertices $v$, $\sigma_{i,j}^{(1)}(v)=0,\partial_{x}\sigma_{i,j}^{(1)}(v)=\partial_{x}\sigma_{i,j}^{(1)}(v)=\frac{1}{2}q_{h}(v),$ and $\partial_{xx}\sigma_{ij}^{(1)}(v)=\partial_{x}q_{h}(v),\partial_{yy}\sigma_{i,j}^{(1)}(v)=\partial_{y}q_{h}(v),\partial_{xy}\sigma_{i,j}^{(1)}(v)=0.$
  2. (2)
    For all $e\in\mathcal{E}$ and $p\in P_{k-6}(T;\mathbb{S}^{2})$, $\int_{e}\sigma^{(1)}n_{e}\cdot p=\int_{e}\tau_{h}n_{e}\cdot p.$ Here $n_{e}$ is the unit normal vector of edge $e$.
  3. (3)
    For all $T\in\mathcal{T}$, for all $p\in P_{k-3}^{0}(T;\mathbb{S}^{2})$, $\int_{T}\sigma^{(1)}:p=\int_{T}\tau_{h}:p.$
It can be checked that at each vertex $v\in\mathcal{V}$, we have $\mathrm{div}\sigma^{(1)}(v)=q_{h}(v),~\nabla\mathrm{div}\sigma^{(1)}(v)=\nabla q_{h}(v).$ Moreover, for $k\geq 7$, it holds that
$\int_{T}\mathrm{div}\sigma^{(1)}\cdot p=\int_{\partial T}\sigma^{(1)}n_{\partial T}\cdot p=\int_{\partial T}\tau n_{\partial T}\cdot p=\int_{T}\mathrm{div}\tau\cdot p=\int_{T}q_{h}\cdot p,~p\in\mathbf{RM}.$
This comes from the fact that $p\cdot n_{e}$ is linear function for all $p\in\mathbf{RM}.$
In the second step, we notice that in each $T$, it holds that $q_{h}-\operatorname{div}\sigma^{(1)}\in\mathring{Q}_{k-1}(T)\cap\mathbf{RM}^{\perp}.$ By Corollary 3.1, there exist $\sigma^{(2)}\in\Sigma,~\sigma^{(2)}|_{T}\in\mathring{\Sigma}_{k}(T)$, such that
$\mathrm{div}\sigma^{(2)}=q_{h}-\mathrm{div}\sigma^{(1)},\text{ and }\|\sigma^{2}\|_{H^{1}}\lesssim\|q_{h}-\operatorname{div}\sigma^{(1)}\|_{L^{2}}.$
Let $\sigma=\sigma^{(1)}+\sigma^{(2)}$ we finish the construction.
To complete the proof, it suffices to estimate $\|\sigma\|_{H^{1}}$. Since $\|\sigma\|_{H^{1}}\leq\|\sigma^{(1)}\|_{H^{1}}+\|\sigma^{(2)}\|_{H^{1}}\lesssim\|\sigma^{(1)}\|_{H^{1}}+\|q_{h}\|_{L^{2}}$, it then suffices to estimate $\|\sigma^{(1)}\|_{H^{1}}.$
Let $\tau_{h}\in\Sigma_{k}$ be the Scott-Zhang interpolant of $\tau$. Then $(\sigma^{(1)}-\tau_{h})|_{T}\in\mathcal{P}_{k}(T;\mathbb{S}^{2})$. For $\Sigma_{k}^{2}$, let $\varphi_{v,0,i}$, $\varphi_{v,1,i}$, $\varphi_{v,2,i}$ be the nodal basis functions associated with function value $(\cdot)(v)$ , gradient value $\nabla(\cdot)(v)$ and Hessian value $\nabla^{2}(\cdot)(v)$, respectively. $\varphi_{e,j}$ is the nodal basis function corresponding to degrees of freedom $\int_{e}(\cdot)p_{j}$ and $\varphi_{T,j}$ for $\int_{T}(\cdot)p_{j}.$ Expend $(\sigma^{(1)}-\tau_{h})$ by the set of basis function of $\Sigma(T)$, we then have
$\|\sigma^{(1)}-\tau_{h}\|_{H^{1}(T)}$ $=\|\sum_{i=0}^{2}\sum_{v\in\mathcal{V}(T)}\nabla^{i}(\sigma^{(1)}-\tau_{h})(v)\varphi_{v,i}+\sum_{j}\sum_{e\in\mathcal{E}(T)}\int_{e}(\sigma^{(1)}-\tau_{h})p_{j}\varphi_{e,j}$
$+\sum_{j}\int_{T}(\sigma^{(1)}-\tau_{h})p_{j}\varphi_{T,j}\|_{H^{1}(T)}.$
By scaling argument,
$\|\sum_{j}\sum_{v\in\mathcal{V}(T)}\nabla^{i}(\sigma^{(1)}-\tau_{h}):p_{j}(v)\varphi_{v,i,j}\|^{2}_{H^{1}(T)}\lesssim\sum_{a\in\mathcal{V}(T)}(h_{T}^{i}|\nabla^{i}\tau_{h}(a)|^{2}+h_{T}^{2}|q_{h}(a)|^{2}+h_{T}^{4}|\nabla q_{h}(a)|^{2});$
$\|\sum_{j}\sum_{e\in\mathcal{E}(T)}\int_{e}(\sigma^{(1)}-\tau_{h})q_{j}\varphi_{e,j}\|_{H^{1}(T)}=\|\sum_{j}\sum_{e\in\mathcal{E}(T)}\int_{e}(\tau-\tau_{h})q_{j}\varphi_{e,j}\|_{H^{1}(T)}\lesssim h_{T}^{-1/2}\|\tau-\tau_{h}\|_{L^{2}(\partial T)};$
$\|\sum_{j}\int_{T}(\sigma^{(1)}-\tau_{h})q_{j}\varphi_{T,j}\|_{H^{1}(T)}\lesssim\|\sum_{j}\int_{T}(\tau-\tau_{h})q_{j}\varphi_{T,j}\|_{H^{1}(T)}\lesssim h_{T}^{-1}\|\tau-\tau_{h}\|_{L^{2}(T)}.$
By applying a scaling argument and the inverse inequality, we obtain
$\|\sigma^{(1)}-\tau_{h}\|_{H^{1}({T})}\lesssim\|\tau\|_{H^{1}(\widetilde{T})}+\|q_{h}\|_{L^{2}(\widetilde{T})},$
where $\tilde{T}$ denotes the union of all elements adjacent to $T$. Consequently, by shape regularity we have
$\|\sigma^{(1)}\|_{H^{1}}\lesssim\|\tau_{h}\|_{H^{1}}+\|\tau\|_{H^{1}}+\|q_{h}\|_{L^{2}}\lesssim\|q_{h}\|_{L^{2}}.$
3.2. The Pair $\Sigma_{k}^{1}\times Q_{k-1}^{0}$: Matrix-version Falk-Neilan pair
In this subsection, we analyze the results for the pair $\Sigma_{k}^{1}\times Q_{k-1}^{0}$, provided that $k\geq 7$. First, we see how the rank deficiency arises for singular vertices. Recall that for each Type I singular vertex $z$, we denote the faces in the patch surrounding $z$ by $T_{1},T_{2},T_{3},T_{4}$ in counterclockwise order.
Now we fix one of its surrounding elements $T$. Recall the conventional notations in Figure 6, $n_{+}$ represents the outer normal vector, while $n_{-}$ represents the inner normal vector. Note that See Figure 7 for an illustration. We then define
(3.9) $J_{z,T}=t_{+}n_{-}^{T}+t_{-}n_{+}^{T}.$
$\boldsymbol{t}_{-}$$\boldsymbol{t}_{z}$$\boldsymbol{t}_{+}$$z$$y_{-}$$y_{+}$$\boldsymbol{n}_{-}$$\boldsymbol{n}_{+}$
$\boldsymbol{t}$$\boldsymbol{n}$
Figure 7. Notations in an element.
Lemma 3.1 (Necessary condition for Type I singularity).
Let $z$ be a Type I singular vertex, with four elements $T_{1},T_{2},T_{3},T_{4}$ sharing $z$, as shown in Figure 8. Then for any continuous, piecewise $C^{2}$, symmetric matrix-valued function $\tau\in C^{0}(\operatorname{star}(z);\mathbb{S}^{2})$, it holds that
(3.10) $J_{z,T_{1}}:\nabla\operatorname{div}\tau|_{T_{1}}(z)+J_{z,T_{2}}:\nabla\operatorname{div}\tau|_{T_{2}}(z)+J_{z,T_{3}}:\nabla\operatorname{div}\tau|_{T_{3}}(z)+J_{z,T_{4}}:\nabla\operatorname{div}\tau|_{T_{4}}(z)=0.$
[…]
Figure 8. A configuration for type I singularity.
Proof.
Without loss of generality, we rotate the edge $\overline{zy_{1}}$ to align with the $x$-axis, and set $z$ as the origin. In this case, $\alpha=0$ and we define $\theta$ as the angle between $\overline{zy_{2}}$ and the $x$-axis. Consider the transformation:
$\begin{pmatrix}\tilde{x}\\ \tilde{y}\end{pmatrix}=\begin{pmatrix}1&0\\ \cos\theta&\sin\theta\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}.$
By direct computation within a single element, we obtain the following results when evaluating at $z$:
$\partial_{x}(\mathrm{div}\tau)_{1}$ $=\partial^{2}_{\tilde{x}}\tau_{11}+2\cos\theta\partial^{2}_{\tilde{x}\tilde{y}}\tau_{11}+\sin\theta\partial^{2}_{\tilde{x}\tilde{y}}\tau_{12}+\cos^{2}\theta\partial^{2}_{\tilde{y}}\tau_{11}+\cos\theta\sin\theta\partial^{2}_{\tilde{y}}\tau_{12},$
$\partial_{y}(\mathrm{div}\tau)_{1}$ $=\sin\theta\partial^{2}_{\tilde{x}\tilde{y}}\tau_{11}+\sin\theta\cos\theta\partial^{2}_{\tilde{y}}\tau_{11}+\sin^{2}\theta\partial^{2}_{\tilde{y}}\tau_{12},$
$\partial_{x}(\mathrm{div}\tau)_{2}$ $=\partial^{2}_{\tilde{x}}\tau_{12}+2\cos\theta\partial^{2}_{\tilde{x}\tilde{y}}\tau_{12}+\sin\theta\partial^{2}_{\tilde{x}\tilde{y}}\tau_{22}+\cos^{2}\theta\partial^{2}_{\tilde{y}}\tau_{12}+\sin\theta\cos\theta\partial^{2}_{\tilde{y}}\tau_{22},$
$\partial_{y}(\mathrm{div}\tau)_{2}$ $=\sin\theta\partial^{2}_{\tilde{x}\tilde{y}}\tau_{12}+\sin\theta\cos\theta\partial^{2}_{\tilde{y}}\tau_{12}+\sin^{2}\theta\partial^{2}_{\tilde{y}}\tau_{22}.$
Eliminating the mixed derivatives, we find the following identity holds at $z$:
$J_{z,T_{1}}$ $:\nabla\mathrm{div}\tau|_{T_{1}}=\sin\theta\partial^{2}_{\tilde{x}}\tau_{11}-\sin\theta\cos^{2}\theta\partial^{2}_{\tilde{y}}\tau_{11}-2\cos\theta\sin^{2}\theta\partial^{2}_{\tilde{y}}\tau_{12}-\sin^{3}\theta\partial^{2}_{\tilde{y}}\tau_{22}.$
Since $\tau$ is continuous, it follows that $\partial^{2}_{\tilde{x}}\tau$ is continuous on $\overline{zy_{1}},\overline{zy_{3}}$ and $\partial^{2}_{\tilde{y}}\tau$ is continuous on $\overline{zy_{2}},\overline{zy_{4}}$. Therefore, for the values at $z$ we have
$J_{z,T_{1}}:\nabla\mathrm{div}\tau|_{T_{1}}+J_{z,T_{2}}:\nabla\mathrm{div}\tau|_{T_{2}}+J_{z,T_{3}}:\nabla\mathrm{div}\tau|_{T_{3}}+J_{z,T_{4}}:\nabla\mathrm{div}\tau|_{T_{4}}=0.$
Remark 3.4.
For the singular case, it holds that $J_{z}:=J_{z,1}=-J_{z,2}=J_{z,3}=-J_{z,4}.$ Therefore, (3.10) can be rewritten as an alternating sum:
$J_{z}:\sum_{s=1}^{4}(-1)^{s}\nabla\operatorname{div}\tau(z)|_{T_{s}}=0.$
Suppose that $\tau\in\Sigma_{k}^{0}$. If we set $q=\operatorname{div}\tau$, then we have
(3.11) $J_{z}:\nabla q_{1}(z)+J_{z}:\nabla q_{3}(z)=J_{z}:\nabla q_{2}(z)+J_{z}:\nabla q_{4}(z).$
Here, $q_{s}=q|_{T_{s}}.$
Now we prove the sufficiency. We need to show that for all multi-valued vector-valued functions $q$ such that (3.11) holds, there exists $\sigma\in\Sigma_{k}^{0}(z)$ such that $\operatorname{div}\sigma(z)=q(z)$. The challenge arises here from the nontrivial null space of the mapping $\tau\mapsto J_{z}:\tau$ in (3.11). To resolve this, we deal with two different types of $\tau$ separately: the first type contributes nothing to both terms in (3.11), and the second type has nonzero but equal contributions to both sides.
Define the null space $\mathcal{N}_{z,T}:=\{\tau\in\mathbb{R}^{2\times 2}:J_{z,T}:\tau=0\}.$ Now we fix an element $T$ and recall the notation in Figure 7. Let
(3.12) $\varphi_{1}=\psi_{z}^{2}\psi_{+}\psi_{-}t_{-}t_{-}^{T},\quad\varphi_{2}=\psi_{z}^{2}\psi_{+}\psi_{-}t_{+}t_{+}^{T},\quad\varphi_{3}=\psi_{z}^{2}\psi_{+}\psi_{-}(t_{-}t_{+}^{T}+t_{+}t_{-}^{T}).$
Clearly, all $\varphi_{i}$ for $i=1,2,3$ are supported in $T$, with $\nabla\varphi_{i}(v)=0$ for all $v\in\mathcal{V}$ and $\nabla^{2}\varphi_{i}(v)=0$ for $v\neq z$. Consequently, $\varphi_{i}\in\Sigma_{4}^{1}(\operatorname{star}(z))$, and vanishes on the boundary. We now evaluate $\nabla\operatorname{div}\varphi_{i}(z)$ to see they form a basis of the null space.
Lemma 3.2.
Fix $z,T$ and let $\varphi_{i}$ be defined as (3.12). Then, $\nabla\operatorname{div}\varphi_{i}(z),i=1,2,3$ form a basis of $\mathcal{N}_{z,T}=\{\tau\in\mathbb{R}^{2\times 2}:J_{z,T}:\tau=0\}.$
Proof.
Note that $\nabla\psi_{-}=\frac{1}{h_{-}}n_{+},\nabla\psi_{+}=-\frac{1}{h_{+}}n_{-}$, and $t_{+}\cdot n_{-}=-t_{-}\cdot n_{+}=\sin\theta$, where $\theta$ is the angle between $t_{+}$ and $t_{-}$. By direct calculation, it holds that
(3.13) $\nabla\operatorname{div}\varphi_{1}(z)$ $=\frac{\sin\theta}{h_{+}h_{-}}n_{-}t_{-}^{T},$
$\nabla\operatorname{div}\varphi_{2}(z)$ $=-\frac{\sin\theta}{h_{+}h_{-}}n_{+}t_{+}^{T},$
$\nabla\operatorname{div}\varphi_{3}(z)$ $=-\frac{\sin\theta}{h_{+}h_{-}}(n_{-}t_{+}^{T}-n_{+}t_{-}^{T}).$
Therefore, it can be checked that $\nabla\operatorname{div}\varphi_{i}$ form a basis.
We now proceed with the proof through a case-by-case analysis, starting with the nonsingular case. The proof strategy is similar to that of GuzmanScott19 . Suppose that $z\in\mathcal{V}\setminus\mathcal{V}_{I}$ is a nonsingular vertex. For any multi-valued vector-valued function $q\in Q_{k-1}^{0}$, we now find $\sigma\in\Sigma_{k}^{1}$ such that $\operatorname{div}\sigma(z)=q(z)$. Furthermore, we obtain bounded estimates whenever $\Theta_{I}(x)$ is bounded below (no nearly singular vertices). The following lemma provides the details.
Lemma 3.3 (Gradient Value Modification for the Nonsingular Case).
For $z\in\mathcal{V}\setminus\mathcal{V}_{I}$ and any $p\in Q_{k-1}^{0}$, there exists $\sigma\in\Sigma_{k}^{1}$ satisfying:
  1. (1)
    $\sigma$ is supported in $\operatorname{star}(z)$.
  2. (2)
    $\mathrm{div}\sigma(v)=0$ for all $v\in\mathcal{V}$.
  3. (3)
    $\nabla\mathrm{div}\sigma(v)=0$ for all $v\neq z$.
  4. (4)
    $\nabla\mathrm{div}\sigma|_{T}(z)=\nabla q|_{T}(z)$ for all $T\in\operatorname{star}(z)$.
Moreover, the bound $\|\sigma\|_{H^{1}}\lesssim\frac{1}{\Theta_{I}(z)^{2}}\|q\|_{L^{2}(\operatorname{star}(z))}$ holds.
Proof.
The construction proceeds in several steps, with the following strategy. We first construct $\tau$ such that $\nabla\operatorname{div}\tau-\nabla q$ lies in the null space $\mathcal{N}_{z,T}$ for each $T\in\operatorname{star}(z)$. Then, we apply Lemma 3.2 to adjust the values lying in the null space. For brevity, we denote $J_{z,s}=J_{z,T_{s}}$ and $\mathcal{N}_{z,s}=\mathcal{N}_{z,T_{s}}$, where $J_{z,s}=t_{s}n_{s-1}^{T}+t_{s-1}n_{s}^{T}.$ Recall that this agrees with (3.9).
Step 1: We first consider $T_{s}$ with $\sin(\theta_{s}+\theta_{s-1})=\Theta_{I}(z)>0$ and show how to modify the gradient value in $T_{s}$. For such $s$, we define $\sigma_{s}^{+}=\psi_{z}^{3}\psi_{s}^{2}t_{s+1}t_{s+1}^{T}$. This $\sigma_{s}^{+}$ satisfies conditions (1), (2), and (3) of Lemma 3.3. For condition (4), we compute the value separately:
  • In $T_{s+1}$, we have
    $\nabla\operatorname{div}\sigma_{s}^{+}|_{T_{s+1}}(z)=(t_{s+1}\cdot\nabla\lambda_{y_{s}})\nabla\lambda_{y_{s}}t_{s+1}^{T}=0.$
  • In $T_{s}$, we have
    $\nabla\operatorname{div}\sigma_{s}^{+}|_{T_{s}}(z)=(\nabla\lambda_{y_{s}}\cdot t_{s+1})\nabla\lambda_{y_{s}}t_{s+1}^{T}=-\frac{\sin(\theta_{s}+\theta_{s+1})}{h_{s}^{2}}n_{s-1}t_{s+1}^{T}.$
Thus,
$J_{z,s}:\nabla\operatorname{div}\sigma_{s}^{+}|_{T_{s}}(z)$ $=(t_{s}n_{s-1}^{T}+t_{s-1}n_{s}^{T}):\left(-\frac{\sin(\theta_{s}+\theta_{s+1})}{h_{s}^{2}}n_{s-1}t_{s+1}^{T}\right)$
$=\frac{\sin\theta_{s}}{h_{s}^{2}}\sin^{2}(\theta_{s}+\theta_{s+1}).$
Since $\sin(\theta_{s}+\theta_{s-1})\neq 0$, we define
$\tau_{s}=\frac{h_{s}^{2}(J_{z,s}:\nabla q(z))}{\sin\theta_{s}\sin^{2}(\theta_{s}+\theta_{s+1})}\sigma_{s}^{+},$
ensuring $J_{z,s}:\nabla\operatorname{div}\tau_{s}(z)=J_{z,s}:\nabla q(z)$ in $T_{s}$. Moreover,
$\|\tau_{s}\|_{H^{1}}\lesssim\frac{h_{s}^{2}|\nabla q(z)|}{\Theta_{I}(z)^{2}}\lesssim\frac{1}{\Theta_{I}(z)^{2}}\|q\|_{L^{2}(\operatorname{star}(z))}.$
Here the last inequality comes from the scaling argument.
Step 2: We now adjust the value for other $T_{s^{\prime}}$, $s^{\prime}\neq s$ to make sure the modified gradient is in the null space. Without loss of generality, assume $s^{\prime}=1$. For each $r$, define $\sigma_{r}=\psi_{z}^{3}\psi_{r}^{2}t_{r}t_{r}^{T}$. In $T_{r}$,
$\nabla\operatorname{div}\sigma_{r}|_{T_{r}}(z)=(t_{r}\cdot\nabla\lambda_{r})\nabla\lambda_{r}t_{r}^{T}=-\frac{\sin\theta_{r}}{h_{r}^{2}}n_{r-1}t_{r}^{T},$
thus it holds that
$J_{z,r}:\nabla\operatorname{div}\sigma_{r}|_{T_{r}}(z)=(t_{r}n_{r-1}^{T}+t_{r-1}n_{r}^{T}):\left(-\frac{\sin\theta_{r}}{h_{r}^{2}}n_{r-1}t_{r}^{T}\right)=-\frac{\sin^{3}\theta_{r}}{h_{r}^{2}}.$
In $T_{r+1}$,
$\nabla\operatorname{div}\sigma_{r}|_{T_{r+1}}(z)=\frac{\sin\theta_{r}}{h_{r}^{2}}n_{r+1}t_{r}^{T},$
and
$J_{z,r+1}:\nabla\operatorname{div}\sigma_{r}|_{T_{r+1}}(z)=\frac{\sin^{3}\theta_{r+1}}{h_{r+1}^{2}}.$
Let $\tilde{\sigma}_{r}=\frac{h_{r}^{2}}{\sin^{3}\theta_{r}}\sigma_{r}$. Then we have
$\|\tilde{\sigma}_{r}\|_{H^{1}}\lesssim h_{r}^{2},\quad|J_{z,r+1}:\nabla\operatorname{div}\tilde{\sigma}_{r}|_{T_{r+1}}(z)|\lesssim 1.$
Define
$\tilde{\tau}_{s}=\alpha_{s-1}\frac{h_{s}^{2}(J_{z,s}:\nabla q(z))}{\sin\theta_{s}\sin^{2}(\theta_{s}+\theta_{s+1})}\sigma_{s}^{+},$
where $\sigma_{s}^{+}$ is defined in Step 1. Now let
$\tau_{1}=\alpha_{1}\tilde{\sigma}_{1}+\cdots+\alpha_{s-1}\tilde{\sigma}_{s-1}+\tilde{\tau}_{s},$
where $\alpha_{1}=J_{z,1}:\nabla q(z)$, and $\alpha_{r}=\alpha_{r-1}(J_{z,r}:\nabla\operatorname{div}\tilde{\sigma}_{r-1}|_{T_{r}}(z))$ for $r=2,\ldots,s-1$, sequentially. Thus, $|\alpha_{r}|\lesssim|\alpha_{1}|\lesssim|\nabla q(z)|$, and
$\|\alpha_{r}\tilde{\sigma}_{r}\|_{H^{1}}\lesssim|\nabla q(z)|h_{r}^{2}\lesssim\|q\|_{L^{2}(\operatorname{star}(z))},\quad\|\tilde{\tau}_{s}\|_{H^{1}}\lesssim\frac{1}{\Theta_{I}(z)^{2}}\|q\|_{L^{2}(\operatorname{star}(z))}.$
This $\tau_{1}$ satisfies $J_{z,1}:\nabla\operatorname{div}\tau_{1}|_{T_{1}}(z)=J_{z,1}:\nabla q(z)$, while $J_{z,s}:\nabla\operatorname{div}\tau_{1}|_{T_{s}}(z)=0$ for $s\neq 1$. Similarly, we can construct $\tau_{s^{\prime}}$ for other $s^{\prime}\neq 1$.
Step 3: After the previous two steps, there exists $\tau\in\Sigma_{0}^{k}$ such that $\nabla\operatorname{div}\tau-\nabla q$ to lie in $\mathcal{N}_{z,T}$ for all $T\in\operatorname{star}(z)$. Using Lemma 3.3, we conclude the proof. ∎
Remark 3.5.
The above argument assumes that $z$ is an interior vertex. For the case that $z$ is a boundary vertex, we can choose any boundary element $T_{s}$ in Step 1 and then continue the remaining proof.
Now we focus on the singular case. The idea behind the proof is similar to the non-singular case.
Lemma 3.4 (Gradient Value Modification for the Singular Case).
For $z\in\mathcal{V}_{I}$, and any $q\in Q_{k-1}^{0}$ such that (3.11) holds. There exists $\sigma\in\Sigma_{k}^{1}$, such that the following condition holds:
  1. (1)
    $\sigma$ is supported in $\operatorname{star}(z)$.
  2. (2)
    $\mathrm{div}\sigma(v)=0,\text{ for all }v\in\mathcal{V}$,
  3. (3)
    $\nabla\mathrm{div}\sigma(v)=0,~\text{for all }v\neq z$.
  4. (4)
    $\nabla\mathrm{div}\sigma|_{T}(z)=\nabla q|_{T}(z)$ for all $T\in\operatorname{star}(z)$.
Moreover, it holds that $\|\sigma\|_{H^{1}}\lesssim\|q\|_{L^{2}(\operatorname{star}(z))}.$
Proof.
The proof is similar to the second step of Lemma 3.3. Note in this case we have $J_{z,r+1}:\nabla\operatorname{div}\tilde{\sigma}_{r}|_{T_{r+1}}(z)=1$. Therefore, we finish the proof. ∎
The following theorem gives the full characterization of $\operatorname{div}\Sigma_{k}^{1}$.
Theorem 3.3.
The divergence of the space $\Sigma_{k}^{1}$ is characterized as
$\operatorname{div}\Sigma_{k}^{1}=\left\{q\in Q_{k-1}^{0}:J_{z}:\nabla q_{1}(v)+J_{z}:\nabla q_{3}(v)=J_{z}:\nabla q_{2}(v)+J_{z}:\nabla q_{4}(v),\ \forall v\in\mathcal{V}_{I}\right\},$
where $q_{i}:=q|_{T_{i}}$ denotes the restriction of $q$ to element $T_{i}$ around a singular vertex. Moreover, if $\Theta_{I}(x)\geq\theta_{I}>0$ for all $x\in\mathcal{V}\setminus\mathcal{V}_{I}$, then for any $q\in\operatorname{div}\Sigma_{k}^{1}$, there exists $\sigma\in\Sigma_{k}^{1}$ such that
$\operatorname{div}\sigma=q\quad\text{and}\quad\|\sigma\|_{H^{1}}\lesssim\theta_{I}^{-2}\|q\|_{L^{2}}.$
Proof.
By Lemmas 3.3 and 3.4, for each vertex $z\in\mathcal{V}$, there exists $\sigma_{z}\in\Sigma_{k}^{1}$ such that:
  • $\sigma_{z}$ is supported in $\operatorname{star}(z)$,
  • $\nabla\operatorname{div}\sigma_{z}(v)=0$ for all $v\neq z$,
  • $\nabla\operatorname{div}\sigma_{z}|_{T}(z)=\nabla q|_{T}(z)$ for all $T\in\operatorname{star}(z)$,
with the bound
$\|\sigma_{z}\|_{H^{1}}\lesssim\theta_{I}^{-2}\|q\|_{L^{2}(\operatorname{star}(z))}.$
Define $\sigma^{(1)}=\sum_{z\in\mathcal{V}}\sigma_{z}$. Then, $\nabla\operatorname{div}\sigma^{(1)}-\nabla q$ vanishes at each vertex in $\mathcal{V}$. Since each $\sigma_{z}$ is supported in $\operatorname{star}(z)$, it follows that
$\|\sigma^{(1)}\|_{H^{1}}\lesssim\theta_{I}^{-2}\|q\|_{L^{2}}.$
Next, observe that $q^{(1)}:=q-\operatorname{div}\sigma^{(1)}\in Q_{k-1}^{1}$. By the surjectivity of the divergence operator from $\Sigma_{k}^{2}$ to $Q_{k-1}^{1}$, there exists $\sigma^{(2)}\in\Sigma_{k}^{2}$ such that
$\operatorname{div}\sigma^{(2)}=q^{(1)}=q-\operatorname{div}\sigma^{(1)},$
with
$\|\sigma^{(2)}\|_{H^{1}}\lesssim\|q^{(1)}\|_{L^{2}}\leq\|q\|_{L^{2}}+\|\operatorname{div}\sigma^{(1)}\|_{L^{2}}\lesssim\theta_{I}^{-2}\|q\|_{L^{2}}.$
Finally, set $\sigma=\sigma^{(1)}+\sigma^{(2)}$. Then,
$\operatorname{div}\sigma=\operatorname{div}\sigma^{(1)}+\operatorname{div}\sigma^{(2)}=\operatorname{div}\sigma^{(1)}+(q-\operatorname{div}\sigma^{(1)})=q,$
and
$\|\sigma\|_{H^{1}}\leq\|\sigma^{(1)}\|_{H^{1}}+\|\sigma^{(2)}\|_{H^{1}}\lesssim\theta_{I}^{-2}\|q\|_{L^{2}},$
completing the proof. ∎
3.3. The pair $\Sigma_{k}^{0}\times Q_{k-1}^{-1}$: Matrix-version Scott–Vogelius pair
This subsection establishes the characterization and stability of the divergence image of the symmetric tensor Lagrange elements, which can be regarded as a matrix-version of the Scott–Vogelius pair. As previously noted, the primary challenge arises from the presence of two types of singularities, with Type II singularities being of particular interest in the stress gradient equation.
We first present the necessary results for both types of singularities.
Lemma 3.5 (Necessary Conditions for Type I Singularity).
Suppose $\sigma$ is continuous and piecewise $C^{2}$, with $q=\operatorname{div}\sigma$. Then, at any vertex $v\in\mathcal{V}_{I}$, the following equations hold:
(3.14) $q_{1}(v)+q_{3}(v)=q_{2}(v)+q_{4}(v),$
(3.15) $J_{z}:\nabla q_{1}(v)+J_{z}:\nabla q_{3}(v)=J_{z}:\nabla q_{2}(v)+J_{z}:\nabla q_{4}(v).$
Proof.
It suffices to prove (3.14), as (3.15) is addressed in (3.11). The condition in (3.14) mirrors the analysis of the Scott-Vogelius Stokes pair. The symmetric tensors $t_{1}t_{1}^{T}$, $t_{2}t_{2}^{T}$, and $t_{1}t_{2}^{T}+t_{2}t_{1}^{T}$ form a basis of $\mathbb{S}^{2}$. Thus, we can express
$\sigma=\alpha t_{1}t_{1}^{T}+\beta t_{2}t_{2}^{T}+\gamma(t_{1}t_{2}^{T}+t_{2}t_{1}^{T}),$
where $\alpha$, $\beta$, and $\gamma$ are continuous Lagrange functions. The divergence is
$\operatorname{div}\sigma|_{T_{i}}(z)=(\partial_{t_{1}}\alpha_{i}+\partial_{t_{2}}\gamma_{i})t_{1}+(\partial_{t_{2}}\beta_{i}+\partial_{t_{1}}\gamma_{i})t_{2},$
where $\alpha_{i}=\alpha|_{T_{i}}(z)$, and $\beta_{i}$, $\gamma_{i}$ are similarly defined. Since $\sigma$ is continuous across elements, evaluating at $v\in\mathcal{V}_{I}$ and summing over elements yields (3.14), completing the proof. ∎
Now we focus on the Type II singularity vertex. We first consider the non-degenerate case. Let $T_{1},\cdots,T_{6}$ be the faces, with vertices $y_{1},\cdots,y_{6}$, and the angle $\theta_{1},\cdots,\theta_{6}$. See Figure 9 for an illustration.
[…]
Figure 9. An illustration for type II singularity, nondegenerate vertex.
With these notations, we have the following lemma, showing the necessary conditions for type II singular vertices.
Lemma 3.6 (Necessary Conditions for Type II Singularity).
Let $z$ be a Type II singular vertex, with four elements $T_{1},T_{2},T_{3},T_{4},T_{5},T_{6}$ sharing $z$, as shown in Figure 9. Then for any continuous, piecewise $C^{1}$, symmetric matrix-valued function $\sigma\in C^{0}(\operatorname{star}(z);\mathbb{S}^{2})$, it holds that
(3.16) $\sum_{i=1}^{6}(-1)^{i}\sin\theta_{i}\,\operatorname{div}\sigma_{i}(z)\cdot n_{i+1}=0,$
where $\sigma_{i}=\sigma|_{T_{i}}$ is the restriction of $\sigma$ to element $T_{i}$.
The detailed proof of Lemma 3.6 is similar to that of Lemma 3.1, and we show the proof in the appendix.
Remark 3.6.
To see how (3.16) is derived, we can let $\sigma$ be the Lagrange $P_{1}$ finite element space $\lambda_{s}\tau$, and $\tau$ be any piecewise constant symmetric matrix. We then have
$\operatorname{div}\sigma_{s}(z)=\frac{\tau\cdot n_{s-1}}{|zy_{s}|\sin\theta_{s}}\text{ and }\operatorname{div}\sigma_{s+1}(z)=-\frac{\tau\cdot n_{s+1}}{|zy_{s}|\sin\theta_{s+1}}.$
Since $n_{s-1}\cdot n_{s+1}=-n_{s+1}\cdot n_{s+2}$, we can see that any $\sigma=\lambda_{s}\tau$ satisfies (3.16).
Remark 3.7 (Necessary case on degenerate case).
We make a remark here on the degenerate case. A straightforward observation is that to obtain the corresponding necessary condition, we can pretend the vertex is non-degenerate by elongating each line and then write down the corresponding equation (3.16). Therefore, we adopt the same notation as in the non-degenerate case. Namely, we use $T_{i}$, $t_{i}$, $n_{i}$ for $i=1,2,\cdots,6$ for the imaging elements and vectors, and use the notations $T_{s,s+1}$ to denote the actual elements. Therefore, we adhere to the unified condition (3.16). See Figure 9 for an illustration. Note that this actually leaves lots of technical details to the verification of the sufficiency, see Appendix A.
[…]
Figure 10. An illustration of Type II singularities: degenerate vertices (Type II-1, II-2, II-3 from left to right).
Now we consider the sufficiency part. The following decomposition will be frequently used:
$u=\frac{u\cdot n_{s}}{n_{s}\cdot t_{s-1}}t_{s-1}+\frac{u\cdot n_{s-1}}{n_{s-1}\cdot t_{s}}t_{s}.$
We first focus on the non-singular vertex, which is illustrated in Figure 11.
[…]
Figure 11. An illustration for $T_{s}$ for a non-singular vertex.
Lemma 3.7 (Sufficient Conditions for Nonsingular Vertex).
Suppose $z\in\mathcal{V}\setminus(\mathcal{V}_{I}\cup\mathcal{V}_{II})$. Then, for any $q\in Q_{k-1}^{0}$, there exists $\sigma\in\Sigma_{k}^{0}$ such that:
  1. (1)
    $\sigma$ is supported in $\operatorname{star}(z)$.
  2. (2)
    $\operatorname{div}\sigma(v)=0$ for all $v\neq z$.
  3. (3)
    $\operatorname{div}\sigma|_{T}(z)=q|_{T}(z)$ for all $T\in\operatorname{star}(z)$.
Moreover, the bound holds:
$\|\sigma\|_{H^{1}}\lesssim\Theta_{II}(z)^{-2}\|q\|_{L^{2}(\operatorname{star}(z))}.$
Proof.
The construction proceeds in several steps.
Step 1: Choose an element $T_{s}\in\operatorname{star}(z)$ such that $|\sin(\theta_{s}+\theta_{s-1})|\geq\Theta_{II}(z)>0$ and $|\sin(\theta_{s}+\theta_{s+1})|\geq\Theta_{II}(z)>0$. Define
$\varphi_{s,-}=\psi_{s-1}\psi_{z}^{2}t_{s-2}t_{s-2}^{T},\quad\varphi_{s,+}=\psi_{s}\psi_{z}^{2}t_{s+1}t_{s+1}^{T}.$
In $T_{s}$, we have:
$\operatorname{div}\varphi_{s,-}|_{T_{s}}(z)=\frac{\sin(\theta_{s}+\theta_{s-1})}{h_{s-2}}t_{s-2},\quad\operatorname{div}\varphi_{s,+}|_{T_{s}}(z)=\frac{\sin(\theta_{s}+\theta_{s+1})}{h_{s}}t_{s+1}.$
Since $\sin(\theta_{s}+\theta_{s\pm 1})\neq 0$ and $t_{s-2}\neq\pm t_{s+1}$, the vectors $\operatorname{div}\varphi_{s,-}|_{T_{s}}(z)$ and $\operatorname{div}\varphi_{s,+}|_{T_{s}}(z)$ form a basis for $\mathbb{R}^{2}$. Thus, for $q_{s}(z)=q|_{T_{s}}(z)$, there exist coefficients $\alpha$ and $\beta$ such that
$\alpha\operatorname{div}\varphi_{s,-}|_{T_{s}}(z)+\beta\operatorname{div}\varphi_{s,+}|_{T_{s}}(z)=q_{s}(z),$
where
$\alpha=\frac{q_{s}(z)\cdot n_{s+1}h_{s-2}}{(t_{s-2}\cdot n_{s+1})\sin(\theta_{s}+\theta_{s-1})},\quad\beta=\frac{q_{s}(z)\cdot n_{s-2}h_{s}}{(t_{s+1}\cdot n_{s-2})\sin(\theta_{s}+\theta_{s+1})}.$
By shape regularity, $|\alpha|,|\beta|\lesssim\frac{|q_{s}(z)|h}{\Theta_{II}(z)^{2}}$. Next, we
$\tau_{s}=\alpha\varphi_{s,-}+\beta\varphi_{s,+},$
so
$\|\tau_{s}\|_{H^{1}}\lesssim\frac{|q_{s}(z)|h}{\Theta_{II}(z)^{2}}\lesssim\Theta_{II}(z)^{-2}\|q\|_{L^{2}(\operatorname{star}(z))}.$
Thus, $\tau_{s}$ satisfies:
  1. (1)
    $\operatorname{div}\tau_{s}|_{T}(z)=0$ for $T\neq T_{s}$,
  2. (2)
    $\operatorname{div}\tau_{s}|_{T_{s}}(z)=q_{s}(z)$,
  3. (3)
    $\operatorname{div}\tau_{s}(y)=0$ for all vertices $y\neq z$,
  4. (4)
    $\|\tau_{s}\|_{H^{1}}\lesssim\Theta_{II}(z)^{-2}\|q\|_{L^{2}(\operatorname{star}(z))}$.
Step 2: Without loss of generality, let $s^{\prime}=0$. For each $r$ and $u\in\mathbb{R}^{2}$, define
(3.17) $\xi_{r}(u)=\psi_{r}\psi_{z}^{2}\frac{u\cdot n_{r-1}}{n_{r-1}\cdot t_{r}}\frac{h_{r}}{\sin\theta_{r}}t_{r}t_{r}^{T}+\psi_{r}\psi_{z}^{2}\frac{u\cdot n_{r}}{n_{r}\cdot t_{r-1}}\frac{h_{r-1}}{\sin\theta_{r}}(t_{r}t_{r-1}^{T}+t_{r-1}t_{r}^{T}).$
By definition, $\xi_{r}(u)$ is supported in $T_{r}\cup T_{r+1}$. In $T_{r}$,
$\operatorname{div}\left(\psi_{r}\psi_{z}^{2}t_{r}t_{r}^{T}\right)|_{T_{r}}(z)=\frac{\sin\theta_{r}}{h_{r}}t_{r},\quad\operatorname{div}\left(\psi_{r}\psi_{z}^{2}(t_{r}t_{r-1}^{T}+t_{r-1}t_{r}^{T})\right)|_{T_{r}}(z)=\frac{\sin\theta_{r}}{h_{r-1}}t_{r-1},$
so
(3.18) $\operatorname{div}\xi_{r}(u)|_{T_{r}}(z)=u.$
In $T_{r+1}$,
$\operatorname{div}\left(\psi_{r}\psi_{z}^{2}t_{r}t_{r}^{T}\right)|_{T_{r+1}}(z)=\frac{\sin\theta_{r+1}}{h_{r+1}}t_{r},$
$\operatorname{div}\left(\psi_{r}\psi_{z}^{2}(t_{r}t_{r-1}^{T}+t_{r-1}t_{r}^{T})\right)|_{T_{r+1}}(z)=\frac{\sin\theta_{r+1}}{h_{r+1}}t_{r-1}+\frac{\sin(\theta_{r}+\theta_{r+1})}{h_{r+1}}t_{r}.$
Thus,
(3.19) $\|\xi_{r}(u)\|_{H^{1}}\lesssim\frac{|u|h}{\Theta_{II}(z)^{2}},\quad|\operatorname{div}\xi_{r}(u)|_{T_{r+1}}(z)|\lesssim|u|.$
We now define the tensor sequentially:
$\tau^{\prime}_{0}=\xi_{0}(q_{0}(z)),\quad\tau^{\prime}_{1}=\xi_{1}(-\operatorname{div}\tau^{\prime}_{0}|_{T_{1}}(z)),\quad\tau^{\prime}_{r}=\xi_{r}(-\operatorname{div}\tau^{\prime}_{r-1}|_{T_{r}}(z)),\quad r=2,\ldots,s-1.$
In $T_{r}$,
$\operatorname{div}\tau^{\prime}_{r-1}|_{T_{r}}(z)+\operatorname{div}\tau^{\prime}_{r}|_{T_{r}}(z)=0.$
By shape regularity,
$|\operatorname{div}\tau^{\prime}_{r-1}|_{T_{r}}(z)|\lesssim|\operatorname{div}\tau^{\prime}_{r-2}|_{T_{r-1}}(z)|\lesssim\cdots\lesssim|q_{0}(z)|,$
yielding
$\|\tau^{\prime}_{r}\|_{H^{1}}\lesssim\Theta_{II}(z)^{-2}\|q\|_{L^{2}(\operatorname{star}(z))}.$
Define
$\tau_{0}=\tau^{\prime}_{0}+\cdots+\tau^{\prime}_{s-1}+\tau_{s}(-\operatorname{div}\tau^{\prime}_{s-1}|_{T_{s}}(z)),$
which satisfies:
  1. (1)
    $\operatorname{div}\tau_{0}|_{T}(z)=0$ for $T\neq T_{0}$,
  2. (2)
    $\operatorname{div}\tau_{0}|_{T_{0}}(z)=q_{0}(z),$
  3. (3)
    $\operatorname{div}\tau_{0}(y)=0$ for all vertices $y\neq z$,
  4. (4)
    $\|\tau_{0}\|_{H^{1}}\lesssim\Theta_{II}(z)^{-2}\|q\|_{L^{2}(\operatorname{star}(z))}$.
Finally, we can construct $\tau_{s^{\prime}}$ for all other $T_{s^{\prime}}\in\operatorname{star}(z)$, $s^{\prime}\neq s$. Define $\sigma=\sum_{s^{\prime}}\tau_{s^{\prime}}$, which completes the proof.
In what follows, we consider the case of the two types of singularities, which are critical to the overall analysis. We proceed by examining each type of singularity separately. The following lemma focuses on Type I singularities, indicating that (3.14) is a sufficient condition.
Lemma 3.8 (Sufficient Conditions for Type I Singularity).
Suppose $z\in\mathcal{V}_{I}$ and $q\in Q_{k-1}^{0}$ satisfies (3.14). Then, there exists $\sigma\in\Sigma_{k}^{0}$ such that:
  1. (1)
    $\sigma$ is supported in $\operatorname{star}(z)$.
  2. (2)
    $\operatorname{div}\sigma(v)=0$ for all $v\neq z$.
  3. (3)
    $\operatorname{div}\sigma|_{T}(z)=q|_{T}(z)$ for all $T\in\operatorname{star}(z)$.
Moreover, the bound holds:
$\|\sigma\|_{H^{1}}\lesssim\|q\|_{L^{2}(\operatorname{star}(z))}.$
Proof.
From Step 2 of Lemma 3.7, recall the construction of $\xi_{r}(u)$ defined in (3.17). For $z\in\mathcal{V}_{I}$, the geometric configuration implies four elements $\sin\theta_{r+2}=\sin\theta_{r}$ and $\sin(\theta_{r}+\theta_{r+1})=0$. Thus, it holds that
$\operatorname{div}\left(\psi_{r}\psi_{z}^{2}(t_{r}t_{r-1}^{T}+t_{r-1}t_{r}^{T})\right)|_{T_{r+1}}(z)=\frac{\sin\theta_{r+1}}{h_{r+1}}t_{r-1}.$
Since $\operatorname{div}\left(\psi_{r}\psi_{z}^{2}t_{r}t_{r}^{T}\right)|_{T_{r+1}}(z)=\frac{\sin\theta_{r+1}}{h_{r+1}}t_{r}$, the divergence of $\xi_{r}(u)$ in $T_{r+1}$ becomes:
(3.20) $\operatorname{div}\xi_{r}(u)|_{T_{r+1}}(z)=u.$
It then follows from (3.17) that $\operatorname{div}\xi_{r}(u)|_{T_{r}}(z)=u$. Now we define:
$\tau_{1}:=\xi_{1}(q_{1}(z)),\quad\tau_{2}:=\xi_{2}(q_{2}(z)-q_{1}(z)),\quad\tau_{3}:=\xi_{3}(q_{3}(z)-q_{2}(z)+q_{1}(z)).$
Since (3.14) gives $q_{1}(z)+q_{3}(z)=q_{2}(z)+q_{4}(z)$, it follows that $q_{3}(z)-q_{2}(z)+q_{1}(z)=q_{4}(z)$. Thus, $\tau_{3}=\xi_{3}(q_{4}(z))$. Set $\sigma=\tau_{1}+\tau_{2}+\tau_{3}$. Then:
  • In $T_{1}$: $\operatorname{div}\sigma|_{T_{1}}(z)=\operatorname{div}\tau_{1}|_{T_{1}}(z)=q_{1}(z)$, since $\tau_{2},\tau_{3}$ are supported in $T_{2}\cup T_{3}$ and $T_{3}\cup T_{4}$, respectively.
  • In $T_{2}$: $\operatorname{div}\sigma|_{T_{2}}(z)=\operatorname{div}\tau_{1}|_{T_{2}}(z)+\operatorname{div}\tau_{2}|_{T_{2}}(z)=q_{1}(z)+(q_{2}(z)-q_{1}(z))=q_{2}(z)$.
  • In $T_{3}$: $\operatorname{div}\sigma|_{T_{3}}(z)=\operatorname{div}\tau_{2}|_{T_{3}}(z)+\operatorname{div}\tau_{3}|_{T_{3}}(z)=(q_{2}(z)-q_{1}(z))+q_{4}(z)=q_{3}(z)$, using (3.14).
  • In $T_{4}$: $\operatorname{div}\sigma|_{T_{4}}(z)=\operatorname{div}\tau_{3}|_{T_{4}}(z)=q_{4}(z)$.
Since each $\xi_{r}$ is supported in $T_{r}\cup T_{r+1}$ and $\operatorname{div}\xi_{r}(y)=0$ for other vertices $y\neq z$, $\sigma$ satisfies conditions (1) and (2). Finally, note that $\|\xi_{r}(u)\|_{H^{1}}\lesssim|u|h$ (as $\sin\theta_{r}$ is bounded away from zero). Thus:
$\|\tau_{r}\|_{H^{1}}\lesssim\|q\|_{L^{2}(\operatorname{star}(z))},\quad\|\sigma\|_{H^{1}}\leq\sum_{r=1}^{3}\|\tau_{r}\|_{H^{1}}\lesssim\|q\|_{L^{2}(\operatorname{star}(z))},$
completing the proof. ∎
For Type II singularities, we adopt a similar approach to that used for Type I singularities. However, the analysis of Type II singularities introduces additional complexity due to their distinct geometric configurations, which are classified into nondegenerate and degenerate cases. Each configuration requires careful consideration of the local mesh geometry. The following lemma addresses the nondegenerate Type II singularity case.
Lemma 3.9 (Sufficient Conditions for Nondegenerate Type II Singularity).
Suppose $z\in\mathcal{V}_{II}$ is a nondegenerate vertex and $q\in Q_{k-1}^{0}$ satisfies (3.16). Then, there exists $\sigma\in\Sigma_{k}^{0}$ such that:
  1. (1)
    $\sigma$ is supported in $\operatorname{star}(z)$.
  2. (2)
    $\operatorname{div}\sigma|_{T}(z)=q|_{T}(z)$ for all $T\in\operatorname{star}(z)$.
  3. (3)
    $\operatorname{div}\sigma|_{T}(y)=0$ for all $y\neq z$, $T\in\operatorname{star}(z)$.
Moreover, the bound holds:
$\|\sigma\|_{H^{1}}\lesssim\|q\|_{L^{2}(\operatorname{star}(z))}.$
Proof.
The construction is decomposed into two steps. First, we prove that there exists $\sigma\in\Sigma_{k}^{0}$ such that in each $T_{i}$, it holds that
(3.21) $(\operatorname{div}\sigma|_{T_{i}}(z)-q_{i}(z))\cdot n_{i+1}=0.$
Consider $\eta_{s}=\psi_{s}\psi_{z}^{2}t_{s}t_{s}^{T}$. Note that $\eta_{s}$ is supported in $T_{s}\cup T_{s+1}$. In $T_{s}$, we have
$\operatorname{div}\eta_{s}|_{T_{s}}\cdot n_{s+1}=\frac{\sin\theta_{s}\sin\theta_{s+1}}{h_{s}}=\frac{\sin\theta_{s+1}}{|\overline{zy_{s}}|},$
while in $T_{s+1}$,
$\operatorname{div}\eta_{s}|_{T_{s+1}}\cdot n_{s+2}=\frac{\sin\theta_{s}}{|\overline{zy_{s}}|},$
Therefore, by setting $\tau_{s}=\frac{|\overline{zy_{s}}|}{\sin\theta_{s}\sin\theta_{s+1}}\eta_{s}$, we have
$\operatorname{div}\tau_{s}|_{T_{s}}\cdot n_{s+1}\sin\theta_{s}=\operatorname{div}\tau_{s}|_{T_{s+1}}\cdot n_{s+2}\sin\theta_{s+1}=1.$
This is illustrated in Figure 12. Here the number inside each $T_{s}$ records the value $\operatorname{div}\tau|_{T_{s}}\cdot n_{s+1}\sin\theta_{s}$.
111111
111111
111111
$\cdots$
Figure 12. An illustration supporting the proof in the first step.
Define
$\sigma_{1}=(q_{1}(z)\cdot n_{2})\tau_{1},$
$\sigma_{r}=(q_{r}(z)\cdot n_{r+1}-\operatorname{div}\sigma_{r-1}|_{T_{r}}\cdot n_{r+2})\tau_{r},\text{ for }r=2,3,4,5.$
Consider $\sigma=\sigma_{1}+\cdots+\sigma_{5}.$ Then it holds that for $i=1,2,3,4,5$,
$(\operatorname{div}\sigma|_{T_{i}}(z)-q_{i}(z))\cdot n_{i+1}=0.$
Thanks for (3.16), the above equation also holds for $i=6$. Thus, we finish the first step of construction.
For the second step, we construct $\tilde{\sigma}$ such that $(\operatorname{div}\sigma|_{T_{i}}(z)+\operatorname{div}\tilde{\sigma}|_{T_{i}}(z)-q_{i}(z))=0.$ Thanks to the first step, we already have $\operatorname{div}\sigma|_{T_{i}}(z)-q_{i}(z)=\alpha_{i}t_{i+1}$ for some constant $\alpha_{i}$. Now consider $\tilde{\sigma}_{s}=\psi_{s}\psi_{z}^{2}t_{s+1}t_{s+1}^{T}$, which is supported in $T_{s}\cup T_{s+1}$. Then, direct computation yields $\operatorname{div}\tilde{\sigma}_{s}|_{T_{s+1}}=0$, while $\operatorname{div}\tilde{\sigma}_{s}|_{T_{s}}=\frac{\sin\theta_{s-1}}{h_{s}}t_{s+1}$. Therefore, taking $\tilde{\sigma}=\sum_{i=1}^{6}\frac{\alpha_{s}h_{s}}{\sin\theta_{s-1}}\tilde{\sigma}_{s}$ we finish the proof for the nondegenerate case. ∎
The degenerate cases are similar, and the complete proof are displayed in Appendix A.
Combining all the above results, we have the following theorem.
Theorem 3.4.
It holds that
(3.22) $\begin{split}\operatorname{div}\Sigma_{k}^{0}=\Big{\{}&q\in Q_{k-1}^{-1}:\\ &J_{z}:\nabla q_{1}(v)+J_{z}:\nabla q_{3}(v)=J_{z}:\nabla q_{2}(v)+J_{z}:\nabla q_{4}(v),\,\,\forall v\in\mathcal{V}_{I}\\ &q_{1}(v)+q_{3}(v)=q_{2}(v)+q_{4}(v),\,\,\forall v\in\mathcal{V}_{I}\\ &\sum_{i=1}^{6}(-1)^{i}\sin\theta_{i}q_{i}(v)\cdot n_{i+1}=0,\,\,\forall v\in\mathcal{V}_{II}.\Big{\}}\end{split}$
Moreover, if for all $x\in\mathcal{V}\setminus\mathcal{V}_{I}\setminus\mathcal{V}_{II}$, $\Theta_{II}(z)$ has a uniform lower bound $\theta_{II}>0$, and $\mathcal{T}$ is shape regular. Then for each $q\in\operatorname{div}\Sigma_{k}^{0}$, there exists $\sigma\in\Sigma_{k}^{0}$ such that $\operatorname{div}\sigma=q$ and $\|\sigma\|_{H^{1}}\lesssim\theta_{II}^{-2}\|q\|_{L^{2}}$.
Proof.
The proof is similar to Theorem 3.3. ∎
3.4. Error Estimates
Based on the inf-sup stability conditions, the following error estimates are standard. Note that the constant for inf-sup condition and coercivity are independent of both parameters $\lambda$ and $\iota$. Therefore, the error estimates are also parameter robust (3.1).
Theorem 3.5.
Let $(\sigma,u)\in\Sigma\times Q$ be the exact solution of problem (1.2). Let $\Sigma_{h}\times Q_{h}$ be one of the following pairs:
  1. (1)
    $\Sigma_{k}^{2}\times Q_{k}^{1}$, provided $k\geq 7$;
  2. (2)
    $\Sigma_{k}^{1}\times Q_{k}^{0}$, provided $k\geq 7$ and $\Theta_{I}(z)>\theta_{I}$ for all interior vertices and $\theta_{I}>0$;
  3. (3)
    $\Sigma_{k}^{0}\times Q_{k}^{-1}$, provided $k\geq 7$ and $\Theta_{I}(z)>\theta_{I}$ and $\Theta_{II}(z)>\theta_{II}$ for all interior vertices and $\theta_{I},\theta_{II}>0$.
Let $(\sigma_{h},u_{h})\in\Sigma_{h}\times Q_{h}$ be the finite element solution of (3.1). Then
(3.23) $\|\sigma-\sigma_{h}\|_{\iota}+\|u-u_{h}\|_{L^{2}}\lesssim\inf_{\tau_{h}\in\Sigma_{h},~v_{h}\in Q_{h}}(\|\sigma-\tau_{h}\|_{\iota}+\|u-u_{h}\|_{L^{2}}).$
Here the hidden constant is independent with $\iota$ and $\lambda$.
As a corollary, suppose that $\sigma\in H^{k+1}(\Omega;\mathbb{S}^{2})$ and $u\in H^{k}(\Omega;\mathbb{R}^{2})$, then we have
(3.24) $\|\sigma-\sigma_{h}\|_{\iota}+\|u-u_{h}\|_{L^{2}}\lesssim h^{k}(\|\sigma\|_{H^{k+1}}+\|u\|_{H^{k}}).$
Note that the norm of $\sigma$ and $u$ might depend on $\iota$.
3.5. The smoothed stress complex with spline spaces
Denote by the superspline space $U_{k+2}^{3}$ and $U_{k+2}^{2}$ as
$U_{k+2}^{r}:=\{u\in C^{2}(\Omega):u|_{T}\in P_{k+2},u\in C^{r}(\mathcal{V})\}.$
Therefore, the following space is closed.
$0~{\hookrightarrow}~U^{r+2}_{k+2}(\mathcal{T})\xrightarrow{\operatorname{curl}\operatorname{curl}^{T}}\Sigma_{k}^{r}(\mathcal{T})\xrightarrow{\operatorname{div}}Q_{k-1}^{r-1}(\mathcal{T})\to 0$
When $r=2$, the space $U_{k+2}^{4}$ admits a finite element construction. It can be readily checked with the help of bubble complexes, that the cohomology of (3.5) is isomorphic to $\mathbf{P}_{1}\otimes\mathcal{H}_{dR}(\Omega)$, see ChristiansenHu2023 ; HuLiangLin23 ; HuLinZhang25 . However, when $r=0$ or $r=1$, the space $U_{k+2}^{2}$ (or $U_{k+2}^{3}$) does not admit a finite element construction HuLinWuYuan25 , which leads to the difficulty identifying the cohomology. The exception is when $\Omega$ is homeomorphic to a disk. In that case, an analysis about the scalar potentials can make sure that the following sequence is exact.
(3.25) $\mathbf{P}_{1}~{\hookrightarrow}~U^{r+2}_{k+2}(\mathcal{T})\xrightarrow{\operatorname{curl}\operatorname{curl}^{T}}\Sigma_{k}^{r}(\mathcal{T})\xrightarrow{\operatorname{div}}\operatorname{div}\Sigma_{k}^{r}(\mathcal{T})\to 0.$
Similar exactness argument can be established for the Scott-Vogelius elements with applications, cf. GuzmanScott18 .
Theorem 3.6.
When $\Omega$ is topologically trivial, then (3.25) is exact for $r=0,k\geq 5$, $r=1,k\geq 7$ and $r=2,k\geq 7$.
Proof.
It can be checked that if $\sigma\in\Sigma_{k}^{r}(\mathcal{T})$ and $\operatorname{div}\sigma=0$, then there exists $\psi\in U_{k+2}^{r+2}(\mathcal{T})$ such that $\operatorname{curl}\operatorname{curl}^{T}\psi=\sigma.$ It suffices to check the dimension count. By Alfeld-Schumaker’s formula Alfeld1987 , it holds that for $k\geq 5$, we have
(3.26) $\begin{split}\dim U_{k+2}^{2}(\mathcal{T})=&\binom{k+4}{2}+\binom{k+1}{2}|\mathcal{E}^{\circ}|-\big{[}\binom{k+4}{2}-6\big{]}|\mathcal{V}^{\circ}|+3|\mathcal{V}_{I}|+|\mathcal{V}_{II}|\\ =&6|\mathcal{V}|+(3k-6)|\mathcal{E}|+\binom{k+1}{2}|\mathcal{F}|+3|\mathcal{V}_{I}|+|\mathcal{V}_{II}|.\end{split}$
Here we use $|\mathcal{V}|-|\mathcal{E}|+|\mathcal{F}|=1$, $|\mathcal{V}^{\partial}|=|\mathcal{E}^{\partial}|$, and $3|\mathcal{F}|=|\mathcal{E}|+|\mathcal{E}^{\circ}|$. Therefore, we have
(3.27) $\begin{split}&\dim U_{k+2}^{2}(\mathcal{T})-\dim\Sigma_{k}^{0}(\mathcal{T})+\dim\operatorname{div}\Sigma_{k}^{0}(\mathcal{T})\\ =&\dim U_{k+2}^{2}(\mathcal{T})-\dim\Sigma_{k}^{0}(\mathcal{T})+\dim Q_{k-1}^{-1}(\mathcal{T})+3|\mathcal{V}_{I}|+|\mathcal{V}_{II}|\\ =&3(|\mathcal{V}|-|\mathcal{E}|+|\mathcal{F}|)=3.\end{split}$
There are also some results about the dimension formula in the superspline space $U_{k+2}^{3}(\mathcal{T})$. In toshniwal2023 we have
(3.28) $\begin{split}\dim U_{k+2}^{3}=&\binom{k+4}{2}+[(k-1)^{2}-\binom{k-2}{2}]|\mathcal{E}^{\circ}|-[\binom{k+4}{2}-10]|\mathcal{V}^{\circ}|+|\mathcal{V}_{I}|\\ =&10|\mathcal{V}|+(3k-12)|\mathcal{E}|+[\binom{k-5}{2}-3]|\mathcal{F}|+|\mathcal{V}_{I}|\end{split}$
for $k$ sufficiently large. Therefore, we have:
(3.29) $\dim U_{k+2}^{3}(\mathcal{T})-\dim\Sigma_{k}^{1}(\mathcal{T})+\dim\operatorname{div}\Sigma_{k}^{1}(\mathcal{T})=3.$
On the other side, the above calculation shows that the dimension formula of $\dim U_{k+2}^{3}(\mathcal{T})$ is valid when $k\geq 7$.
The general cohomology and machinery of the complexes with splines are beyond the scope of this paper.
4. Numerical Experiments
In this section, we provide some numerical examples for verification and further investigation. We start from an exact solution of $(\sigma_{0},u_{0})$, corresponds to the case when $\iota=0$. The right-hand side will be constructed by $f=\operatorname{div}\sigma_{0}$.
For practical concerns, we pay more attention on the Hermite case $\Sigma_{k}^{1}\times Q_{k-1}^{0}$ and Lagrange case $\Sigma_{k}^{0}\times Q_{k-1}^{-1}$ for $k=4$, whose polynomial degree is lower than assumed in the previous theorems.
4.1. Verification
We take an exact solution $u_{0}$ for problem (2.6) with sufficiently boundary vanishing conditions. Specifically, we choose $u_{0}=\operatorname{curl}(\sin^{3}\pi x\sin^{3}\pi y)+(1/\lambda)x^{3}(1-x)^{3}y^{3}(1-y)^{3}.$ Note that $u_{0}\in H_{0}^{2}([0,1]^{2};\mathbb{R}^{2})$, therefore for any divergence-free matrix-valued functions $\tau$, we have
$(\nabla\mathcal{A}\sigma(u_{0}),\nabla\tau)=(\nabla\mathrm{symgrad}(u_{0}),\nabla\tau)=(u_{0},\operatorname{div}\Delta\tau)=0.$
By (2.7), we have $\sigma_{\iota}=\sigma_{0}:=\mathcal{A}^{-1}\mathrm{symgrad}(u_{0}).$ However, the solution of $u_{\iota}$ do not have an analytic expression. Therefore, we use computational results in $\Sigma_{5}^{2}\times Q_{4}^{1}$ and hsize=1/128=1/128= 1 / 128 as reference solutions. Especially when $\iota=0$, we compute the error by the exact solution (available only when $\iota=0$). We fix $\lambda=10^{5}$, $\mu=0.3$.
To resolve the singularities caused by geometry, we use some specific refinements in mesh generation. For the pair $\Sigma_{k}^{1}\times Q_{k-1}^{0}$, we choose the Hsieh-Clough-Tocher split; while for $\Sigma_{k}^{0}\times Q_{k-1}^{-1}$,we choose Morgan-Scott split mesh, respectively, cf. Figure 13. Hereafter, hsize refers to the mesh size before refinement. As a consequence, the actual size of the mesh would be smaller than hsize.
Figure 13. Meshes refined by Hsieh-Clough-Tocher split and Morgan-Scott split
The error between the reference solution and the numerical solutions is denoted as $\|E_{\sigma}\|_{\iota}$ and $\|E_{u}\|_{L^{2}}$. The numerical results of $\Sigma_{4}^{0}\times Q_{3}^{-1}$ and $\Sigma_{4}^{1}\times Q_{3}^{0}$ is listed in Tables 1 and 2.
Table 1. Numerical experiments of $\Sigma_{4}^{1}\times Q_{3}^{0}$ on Hsich-Clough-Tocher Split
hsize 1/2 1/4 1/6 1/8 1/10 1/12 1/14
$\iota=1$
$\|E_{\sigma}\|_{\iota}$ 2.76e+01 3.68e+00 1.18e+00 4.85e-01 2.30e-01 1.21e-01 6.93e-02
order 2.91 2.80 3.09 3.34 3.51 3.63
$\|E_{u}\|_{L^{2}}$ 9.99e+01 1.03e+01 4.45e+00 2.47e+00 1.43e+00 8.58e-01 5.34e-01
order 3.27 2.08 2.04 2.44 2.81 3.08
$\iota=0.01$
$\|E_{\sigma}\|_{\iota}$ 5.26e+00 5.97e-01 1.32e-01 4.22e-02 1.70e-02 8.05e-03 4.27e-03
order 3.14 3.72 3.96 4.07 4.11 4.12
$\|E_{u}\|_{L^{2}}$ 1.09e+00 1.09e-01 4.59e-02 2.52e-02 1.45e-02 8.64e-03 5.36e-03
order 3.32 2.14 2.09 2.48 2.83 3.10
$\iota=1.0e-4$
$\|E_{\sigma}\|_{\iota}$ 5.14e+00 5.89e-01 1.29e-01 4.09e-02 1.64e-02 7.78e-03 4.13e-03
order 3.13 3.74 4.00 4.08 4.11 4.11
$\|E_{u}\|_{L^{2}}$ 1.24e-01 1.24e-02 2.82e-03 9.49e-04 4.02e-04 1.97e-04 1.08e-04
order 3.32 3.65 3.78 3.85 3.90 3.92
$\iota=1.0e-6$
$\|E_{\sigma}\|_{\iota}$ 5.13e+00 5.87e-01 1.28e-01 4.05e-02 1.62e-02 7.68e-03 4.08e-03
order 3.13 3.75 4.01 4.09 4.11 4.11
$\|E_{u}\|_{L^{2}}$ 1.16e-01 1.19e-02 2.53e-03 7.93e-04 3.17e-04 1.50e-04 8.03e-05
order 3.29 3.81 4.04 4.10 4.09 4.07
$\iota=1.0e-8$
$\|E_{\sigma}\|_{\iota}$ 5.13e+00 5.87e-01 1.28e-01 4.05e-02 1.62e-02 7.68e-03 4.08e-03
order 3.13 3.75 4.01 4.09 4.11 4.11
$\|E_{u}\|_{L^{2}}$ 1.16e-01 1.18e-02 2.53e-03 7.91e-04 3.17e-04 1.50e-04 8.02e-05
order 3.29 3.81 4.04 4.10 4.09 4.07
$\iota=0$
$\|E_{\sigma}\|_{\iota}$ 5.13e+00 5.87e-01 1.28e-01 4.05e-02 1.62e-02 7.68e-03 4.08e-03
order 3.13 3.75 4.01 4.10 4.11 4.11
$\|E_{u}\|_{L^{2}}$ 1.16e-01 1.18e-02 2.53e-03 7.91e-04 3.17e-04 1.50e-04 8.02e-05
order 3.29 3.81 4.04 4.10 4.09 4.07
Table 2. Numerical experiments of $\Sigma_{4}^{0}\times Q_{3}^{-1}$ on Morgan-Scott Split
hsize 1/2 1/4 1/6 1/8 1/10 1/12 1/14
$\iota=1$
$\|E_{\sigma}\|_{\iota}$ 1.92e+01 2.40e+00 6.25e-01 2.31e-01 1.04e-01 5.39e-02 3.10e-02
order 3.00 3.32 3.46 3.56 3.63 3.58
$\|E_{u}\|_{L^{2}}$ 7.79e+01 1.27e+01 4.01e+00 1.63e+00 7.83e-01 4.20e-01 2.45e-01
order 2.62 2.83 3.13 3.29 3.41 3.51
$\iota=0.01$
$\|E_{\sigma}\|_{\iota}$ 3.51e+00 3.37e-01 7.00e-02 2.23e-02 9.13e-03 4.39e-03 2.36e-03
order 3.38 3.88 3.97 4.00 4.02 4.03
$\|E_{u}\|_{L^{2}}$ 8.19e-01 1.28e-01 4.04e-02 1.64e-02 7.85e-03 4.21e-03 2.45e-03
order 2.67 2.85 3.14 3.30 3.42 3.51
$\iota=1.0e-04$
$\|E_{\sigma}\|_{\iota}$ 3.40e+00 3.24e-01 6.76e-02 2.17e-02 8.92e-03 4.30e-03 2.32e-03
order 3.39 3.86 3.95 3.98 4.00 4.01
$\|E_{u}\|_{L^{2}}$ 7.43e-02 6.87e-03 1.46e-03 4.81e-04 2.02e-04 9.92e-05 5.44e-05
order 3.44 3.81 3.87 3.89 3.90 3.91
$\iota=1.0e-06$
$\|E_{\sigma}\|_{\iota}$ 3.39e+00 3.21e-01 6.68e-02 2.14e-02 8.80e-03 4.24e-03 2.29e-03
order 3.40 3.87 3.95 3.98 4.00 4.01
$\|E_{u}\|_{L^{2}}$ 6.88e-02 6.39e-03 1.34e-03 4.32e-04 1.79e-04 8.67e-05 4.70e-05
order 3.43 3.86 3.93 3.95 3.97 3.98
$\iota=1.0e-08$
$\|E_{\sigma}\|_{\iota}$ 3.39e+00 3.21e-01 6.68e-02 2.14e-02 8.80e-03 4.24e-03 2.29e-03
order 3.40 3.87 3.95 3.98 4.00 4.01
$\|E_{u}\|_{L^{2}}$ 6.87e-02 6.39e-03 1.34e-03 4.32e-04 1.79e-04 8.67e-05 4.70e-05
order 3.43 3.86 3.93 3.95 3.97 3.98
$\iota=0$
$\|E_{\sigma}\|_{\iota}$ 3.39e+00 3.21e-01 6.68e-02 2.14e-02 8.80e-03 4.24e-03 2.29e-03
order 3.40 3.87 3.95 3.98 4.00 4.01
$\|E_{u}\|_{L^{2}}$ 6.87e-02 6.39e-03 1.34e-03 4.32e-04 1.79e-04 8.67e-05 4.70e-05
order 3.43 3.86 3.93 3.95 3.93 3.98
4.2. Effect of $\partial_{n}\sigma_{0}$
As suggested by Proposition 2.2, the convergence mode will be different for the case when $\partial_{n}\sigma=0$ and $\partial_{n}\sigma\neq 0$. This is further investigated in this subsection, in particular, $H^{2}$ stability of $\sigma_{h}$, where we choose Lamé constants $\mu=0.3,\lambda=1$ be mild.
To achieve this, we set
$\tilde{u}_{0}=(\sin x\sin y(\sin x-\sin 1)(\sin y-\sin 1),\sin x\sin y(e^{x}-e)(e^{y}-e))^{T}.$
And $\tilde{\sigma}_{0}=\mathcal{A}^{-1}\mathrm{symgrad}(\tilde{u}_{0}).$ Note that for this pair, it holds that $\partial_{n}\tilde{\sigma}_{0}|_{\partial\Omega}\neq 0.$ In contrast, the pair $(\sigma_{0},u_{0})$ discussed in the previous subsection have the consistent boundary condition, namely, $\partial_{n}\sigma_{0}=0$. We test the two right hand sides on 1/2,1/4,,1/1281/2,1/4$\cdots$,1/1281 / 2 , 1 / 4 , ⋯ , 1 / 128 size mesh with the pair $(\Sigma^{2}_{5},Q^{1}_{4}),$ and $\iota$ set as $10^{-4},10^{-5}$ and $10^{-6}$. Figure 14 and 15 show the maximum value of $\sigma_{h,xy}$ on each mesh nodes for two different boundary conditions ($\partial_{n}\sigma_{0}=0$ and $\partial\tilde{\sigma}_{0}\neq 0$), respectively. For $(\sigma_{0},u_{0})$, the computational second-order derivatives in Figure 14 are stable with respect to $\iota$ and $h$, while the discrete $W^{2,\infty}$ stability is not observed in Figure 15.
Current numerical experiments indicate that the optimal convergence rate can be achieved even when $\iota\to 0$, provided the left-hand side is smooth enough. A thorough analysis of the continuous problem and more theoretical explanations for the numerical results are beyond the scope of this paper and will be addressed in future work.
[…]
Figure 14. The error of $\|D^{2}\sigma_{\iota,h}\|_{\infty,h}$ (left) and a profile of $\partial_{xy}\sigma_{\iota,h}$ (right) for $\partial_{n}\sigma_{0}|_{\partial\Omega}=0.$
[…]
Figure 15. The error of $\|\partial_{xy}\tilde{\sigma}_{\iota,h}\|_{\infty,h}$ (left) and a profile of $\partial_{xy}\tilde{\sigma}_{\iota,h}$ (right) for $\partial_{n}\tilde{\sigma}_{0}|_{\partial\Omega}\neq 0.$
5. Concluding Remarks
In this paper, we study the finite element discretization of the linear stress gradient theory. The stability usually relies on the geometric structure. Nevertheless, the following questions still remains open.
  1. (1)
    The error estimates at the current stage are not perfect enough. Notably, the asymptotic behavior of the solution as $\iota\to 0$ is absent due to the lack of higher-order regularity results. For higher-order theory, a boudary layer effects would be investigated. It would be interesting to investigate whether the finite element pairs constructed in this paper exhibit the desired asymptotic behavior.
  2. (2)
    The requirement of the polynomial degree $k\geq 7$ is far from practice concerns. In fact, all the finite element pair considered in this paper can be constructed with smaller polynomial degrees. The reduction of the polynomial degree requirement will be a part of the future work.
  3. (3)
    In this paper we focus on natural boundary condition to simplify the argument while addressing the geometric singularities. For boundary conditions $\sigma n=0$ on $\Gamma_{f}\neq\emptyset$, the dimension of discrete space $\Sigma_{h,\Gamma_{f}}^{r}$ will be affected by the corner points on $\Gamma_{f}$, and the stability result will be affected by the singular vertices on $\Gamma_{f}$. The results can be analyzed similarly as FalkNeilan13 , and the detailed results are beyond the scope of this paper.
Acknowledgement
The authors sincerely thank Prof. Dietmar Gallistl for his guidance in discussions and constructive feedback on the manuscript.
Appendix A Technical Details of Section 3.3
We first show the proof of Lemma 3.6.
Proof of Lemma 3.6.
Since $\sigma$ is continuous, and $\{t_{i}t_{i}^{T}\}_{i=1,2,3}$ forms a basis for $\mathbb{S}^{2}$, we express $\sigma=\alpha t_{1}t_{1}^{T}+\beta t_{2}t_{2}^{T}+\gamma t_{3}t_{3}^{T},$ where $\alpha$, $\beta$, and $\gamma$ are continuous functions. On each element $T_{i}$, we express this as
$\sigma|_{T_{i}}=\alpha_{i}t_{1}t_{1}^{T}+\beta_{i}t_{2}t_{2}^{T}+\gamma_{i}t_{3}t_{3}^{T},$
where $\alpha_{i}=\alpha|_{T_{i}}(z)$, and $\beta_{i}$, $\gamma_{i}$ are similarly defined.
Computing the divergence projected onto $n_{i+1}$, we obtain:
$\operatorname{div}\sigma|_{T_{1}}(z)\cdot n_{2}$ $=(n_{2}\cdot t_{1})\partial_{t_{1}}\alpha_{1}+(n_{2}\cdot t_{3})\partial_{t_{3}}\gamma_{1}=-\sin\theta_{2}\partial_{t_{1}}\alpha_{1}+\sin\theta_{3}\partial_{t_{3}}\gamma_{1},$
$\operatorname{div}\sigma|_{T_{2}}(z)\cdot n_{3}$ $=(n_{3}\cdot t_{2})\partial_{t_{2}}\beta_{2}+(n_{3}\cdot t_{1})\partial_{t_{1}}\alpha_{2}=-\sin\theta_{3}\partial_{t_{2}}\beta_{2}-\sin\theta_{1}\partial_{t_{1}}\alpha_{2},$
$\operatorname{div}\sigma|_{T_{3}}(z)\cdot n_{4}$ $=(n_{4}\cdot t_{3})\partial_{t_{3}}\gamma_{3}+(n_{4}\cdot t_{2})\partial_{t_{2}}\beta_{3}=-\sin\theta_{1}\partial_{t_{3}}\gamma_{3}-\sin\theta_{2}\partial_{t_{2}}\beta_{3},$
$\operatorname{div}\sigma|_{T_{4}}(z)\cdot n_{5}$ $=(n_{5}\cdot t_{1})\partial_{t_{1}}\alpha_{4}+(n_{5}\cdot t_{3})\partial_{t_{3}}\gamma_{4}=\sin\theta_{2}\partial_{t_{1}}\alpha_{4}-\sin\theta_{3}\partial_{t_{3}}\gamma_{4},$
$\operatorname{div}\sigma|_{T_{5}}(z)\cdot n_{6}$ $=(n_{6}\cdot t_{2})\partial_{t_{2}}\beta_{5}+(n_{6}\cdot t_{1})\partial_{t_{1}}\alpha_{5}=\sin\theta_{3}\partial_{t_{2}}\beta_{5}+\sin\theta_{1}\partial_{t_{1}}\alpha_{5},$
$\operatorname{div}\sigma|_{T_{6}}(z)\cdot n_{1}$ $=(n_{1}\cdot t_{3})\partial_{t_{3}}\gamma_{6}+(n_{1}\cdot t_{2})\partial_{t_{2}}\beta_{6}=\sin\theta_{1}\partial_{t_{3}}\gamma_{6}+\sin\theta_{2}\partial_{t_{2}}\beta_{6}.$
Since $\alpha$, $\beta$, and $\gamma$ are continuous, we have $\alpha_{i}=\alpha_{i+1}$, $\beta_{i}=\beta_{i+1}$, and $\gamma_{i}=\gamma_{i+1}$ across adjacent elements. Additionally, $\sin\theta_{i}=\sin\theta_{i+3}$ due to the geometric structure around $z\in\mathcal{V}_{II}$. Summing the terms in (3.16), the contributions of $\partial_{t_{j}}\alpha_{i}$, $\partial_{t_{j}}\beta_{i}$, and $\partial_{t_{j}}\gamma_{i}$ cancel due to the alternating signs and periodicity of $\sin\theta_{i}$, yielding
$\sum_{i=1}^{6}(-1)^{i}\sin\theta_{i}\,q_{i}(z)\cdot n_{i+1}=0,$
as required. ∎
In the rest of this section, we consider the sufficient condition of the degenerate type II singular vertices. This is based on the following identity on imaging meshes:
(1.1) $\sin\theta_{i}n_{i+1}-\sin\theta_{i+1}n_{i+2}=\sin(\theta_{i+1}+\theta_{i})n_{i+3}.$
[…]
Figure 16. An illustruation for type II-1 singularity.
Lemma A.1 (Sufficient Conditions for Type II-1 Singularity).
Suppose $z\in\mathcal{V}_{II}$ is a Type II-1 singularity (see Figure 16), and $q\in Q_{k-1}^{0}$ satisfies (3.16). Then, there exists $\sigma\in\Sigma_{k}^{0}$ such that:
  1. (1)
    $\sigma$ is supported in $\operatorname{star}(z)$.
  2. (2)
    $\operatorname{div}\sigma|_{T}(z)=q|_{T}(z)$ for all $T\in\operatorname{star}(z)$.
  3. (3)
    $\operatorname{div}\sigma|_{T}(y)=0$ for all $y\neq z$, $T\in\operatorname{star}(z)$.
Moreover, the bound holds:
$\|\sigma\|_{H^{1}}\lesssim\|q\|_{L^{2}(\operatorname{star}(z))}.$
Proof.
For a Type II-1 singularity at $z\in\mathcal{V}_{II}$, the star $\operatorname{star}(z)$ consists of five elements $T_{1},T_{2},T_{3},T_{4},T_{5,6}$, where $T_{5,6}$ is a single element sharing edges with $T_{4}$ and $T_{1}$, and $n_{5}=-n_{2}$, $\sin\theta_{4}=\sin\theta_{5,6}$ (see Figure 16). The condition (3.16) simplifies to:
$\sum_{i=1}^{4}(-1)^{i}\sin\theta_{i}q_{i}(z)\cdot n_{i+1}+(-1)^{5}\sin\theta_{5,6}q_{5,6}(z)\cdot n_{2}=0.$
Since $n_{5}=-n_{2}$ and $\sin\theta_{4}=\sin\theta_{5,6}$, this becomes:
(1.2) $\sum_{i=1}^{3}(-1)^{i}\sin\theta_{i}q_{i}(z)\cdot n_{i+1}+\sin\theta_{4}(q_{4}(z)-q_{5,6}(z))\cdot n_{2}=0.$
Recall the function $\xi_{r}(u)$ from (3.17) in Lemma 3.7:
$\xi_{r}(u)=\psi_{r}\psi_{z}^{2}\frac{u\cdot n_{r-1}}{n_{r-1}\cdot t_{r}}\frac{h_{r}}{\sin\theta_{r}}t_{r}t_{r}^{T}+\psi_{r}\psi_{z}^{2}\frac{u\cdot n_{r}}{n_{r}\cdot t_{r-1}}\frac{h_{r-1}}{\sin\theta_{r}}(t_{r}t_{r-1}^{T}+t_{r-1}t_{r}^{T}).$
For Type II-1 singularities, as in Lemma 3.8, the mesh geometry implies $\sin(\theta_{r}+\theta_{r+1})=0$, so (3.20) holds: $\operatorname{div}\xi_{r}(u)|_{T_{r}}(z)=u$ and $\operatorname{div}\xi_{r}(u)|_{T_{r+1}}(z)=u$. For $r=4$, since $T_{5,6}$ is adjacent to $T_{4}$, we have:
$\operatorname{div}\xi_{4}(u)|_{T_{4}}(z)=u,\quad\operatorname{div}\xi_{4}(u)|_{T_{5,6}}(z)=u.$
Define $q^{\prime}=q-\operatorname{div}\xi_{4}(q_{5,6}(z))$. Then:
  • In $T_{5,6}$: $\operatorname{div}\xi_{4}(q_{5,6}(z))|_{T_{5,6}}(z)=q_{5,6}(z)$, so $q^{\prime}_{T_{5,6}}(z)=q_{5,6}(z)-q_{5,6}(z)=0$.
  • In $T_{4}$: $\operatorname{div}\xi_{4}(q_{5,6}(z))|_{T_{4}}(z)=q_{5,6}(z)$, so $q^{\prime}_{T_{4}}(z)=q_{4}(z)-q_{5,6}(z)$.
  • In $T_{i}$, $i=1,2,3$: $\operatorname{div}\xi_{4}(q_{5,6}(z))|_{T_{i}}(z)=0$, so $q^{\prime}_{T_{i}}(z)=q_{i}(z)$.
Thus, $q^{\prime}$ satisfies (3.16) with $q^{\prime}_{T_{5,6}}(z)=0$.
Now, we can apply a construction similar to Lemma 3.8 for the four elements $T_{1},T_{2},T_{3},T_{4}$, which completes the proof. ∎
Using similar techniques, we can prove the results of Type II-2 singularity.
Lemma A.2.
Suppose that $z\in\mathcal{V}_{II}$ is of Type II-2 (see also Figure 17), and $q$ satisfies (3.16). Then there exists $\sigma\in\Sigma_{k}^{0}$ satisfying the conditions in Lemma 3.9.
Proof.
For type II-2, (3.16) reads
$\sin\theta_{1}(q_{5,6}(z)-q_{1}(z))\cdot n_{2}+\sin\theta_{2}(q_{2}(z)-q_{3,4}(z))\cdot n_{3}=0.$
We can consider $q^{\prime}=q-\operatorname{div}\xi_{2}(q_{3,4}(z))-\operatorname{div}\xi_{6}(q_{5,6}(z))$. Clearly, $q^{\prime}$ satisfies (3.16), but $q^{\prime}|_{T_{3,4}}=q^{\prime}|_{T_{5,6}}=0$. The remaining argument is similar to Lemma 3.9 by using $\sigma_{1}$, and $\tilde{\sigma}_{1}$, $\tilde{\sigma}_{2}$.
[…]
Figure 17. Illustration for type II-2 singularity.
[…]
Figure 18. Illustration for type II-3 singularity.
Finally, we consider the Type II-3 singularities.
Lemma A.3.
Suppose that $z\in\mathcal{V}_{II}$ is of Type II-3 (see also Figure 18), and $q$ satisfies (3.16). Then there exists $\sigma\in\Sigma_{k}^{0}$ satisfying the conditions in Lemma 3.9.
Proof.
For type II-3, (3.16) reads
$\sin\theta_{1,2}q_{1,2}(z)\cdot n_{4}+\sin\theta_{3,4}q_{3,4}(z)\cdot n_{6}+\sin\theta_{5,6}q_{5,6}(z)\cdot n_{2}=0.$
Again, we adopt a two-step argument. In the first step, we modify the tangential component. Consider $\eta_{s}=\psi_{s}\psi_{z}^{2}t_{s}t_{s}^{T}$. For $\eta_{2}$, direct calculation yields that
$\operatorname{div}\eta_{2}|_{T_{1,2}}\cdot n_{4}=\frac{\sin\theta_{1,2}\sin\theta_{3,4}}{h_{3}}=\frac{\sin\theta_{3,4}}{|\overline{zy_{2}}|},$
and
$\operatorname{div}\eta_{2}|_{T_{3,4}}\cdot n_{6}=-\frac{\sin\theta_{1,2}\sin\theta_{3,4}}{h_{2}}=-\frac{\sin\theta_{1,2}}{|\overline{zy_{2}}|},$
Therefore, by setting $\tau_{2}=\frac{|\overline{zy_{2}}|}{\sin\theta_{1,2}\sin\theta_{3,4}}\eta_{2}$, we have
$\operatorname{div}\tau_{s}|_{T_{1,2}}\cdot n_{4}\sin\theta_{1,2}=-\operatorname{div}\tau_{s}|_{T_{3,4}}\cdot n_{6}\sin\theta_{3,4}=1.$
Similarly, we can construct $\tau_{4}$, $\tau_{6}$, and an analog argument can ensure that
$(\operatorname{div}\sigma|_{T_{1,2}}(z)-q_{1,2}(z))\cdot n_{4}=(\operatorname{div}\sigma|_{T_{3,4}}(z)-q_{3,4}(z))\cdot n_{6}=(\operatorname{div}\sigma|_{T_{5,6}}(z)-q_{5,6}(z))\cdot n_{2}=0.$
In the second step, we modify the tangential components. Now we suppose $\operatorname{div}\sigma|_{T_{1,2}}(z)-q_{1,2}(z)=\alpha_{1,2}t_{4}.$ Set $\tilde{\sigma}_{1,2}=\psi_{2}\psi_{z}^{2}t_{4}t_{4}^{T}.$ It then holds that $\operatorname{div}\tilde{\sigma}_{1,2}|_{T_{3,4}}=0$, while
$\operatorname{div}\tilde{\sigma}_{1,2}|_{T_{1,2}}=\frac{\sin\theta_{5,6}}{h_{1,2}}t_{4}.$
Similarly, we can construct $\tilde{\sigma}_{3,4}$ and $\tilde{\sigma}_{5,6}$ to finish the proof.